### v1

# Zhuk’s bridges, centralizers, and similarity

Abstract.
This is the second of three papers motivated by the author’s desire to understand and explain “algebraically” one aspect of Dmitriy Zhuk’s proof of the CSP Dichotomy Theorem. In this paper we extend Zhuk’s “bridge” construction to arbitrary meet-irreducible congruences of finite algebras in locally finite varieties with a Taylor term. We then connect bridges to centrality and similarity. In particular, we prove that Zhuk’s bridges and our “proper bridges” (defined in our first paper) convey the same information in locally finite Taylor varieties.
Key words and phrases:
universal algebra, abelian congruence, Taylor variety, bridge, centralizer, similarity
2010 Mathematics Subject Classification:
08A05, 08B99 (Primary), 08A70 (Secondary).
The support of the Natural Sciences and Engineering Research Council (NSERC) of Canada is gratefully acknowledged.
1. Introduction
Arguably the most important result in universal algebra in the last ten years is the positive resolution to the Constraint Satisfaction Problem (CSP) Dichotomy Conjecture, announced independently in 2017 by Andrei Bulatov [2] and Dmitriy Zhuk [17, 18]. One particular feature of Zhuk’s proof is his analysis of “rectangular critical” subdirect products $R\leq_{sd}{\mathbf{A}}_{1}\times\cdots\times{\mathbf{A}}_{n}$ of finite algebras ${\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{n}$ in certain locally finite idempotent Taylor varieties. Zhuk showed that such relations $R$ induce derived relations, which he named “bridges,” between certain meet-irreducible congruences of ${\mathbf{A}}_{1},\ldots,{\mathbf{A}}_{n}$ which $R$ determines. Zhuk also established a number of useful properties of his bridges, and ultimately used them to tease out implicit linear equations in CSP instances. In this and two companion papers [16, 15], we aim to understand “algebraically” Zhuk’s bridges and their application to rectangular critical relations.
Our goal in this paper is to establish precise connections between Zhuk’s bridges, centrality, and a relation called “similarity” due to Freese [4] in the congruence modular setting and extended to varieties with a weak difference term (including locally finite Taylor varieties) in our first paper [16]. In Section 2 we give the basic definitions and tools needed in this paper. In Section 3 we summarize the results about similarity from [16] which we will use here. With these preliminaries out of the way, we address two technical limitations of Zhuk’s original presentation: Zhuk defined his bridges between pairs $({\mathbf{A}},\rho)$ and $({\mathbf{B}},\sigma)$ where (i) ${\mathbf{A}}$ and ${\mathbf{B}}$ are finite algebras in a very special kind of Taylor variety, and (ii) $\rho$ and $\sigma$ are congruences satisfying an “irreducibility” property stronger than meet-irreducibility. In fact, Zhuk’s definitions, and all but one of Zhuk’s theorems about bridges (see Theorem 6.4), work in arbitrary locally finite Taylor varieties, so in Section 4 we present his definitions and basic results avoiding limitation (i). Then in Section 5 we use tame congruence theory to show how Zhuk’s definitions and basic results can extend to arbitrary meet-irreducible congruences, avoiding limitation (ii). Finally, in Section 6 we align Zhuk’s bridges, in this broader context, with the algebraic relations of centrality and similarity. In particular, we show that Zhuk’s bridges and our “proper bridges” defined in [16] (see Definition 3.7) are equivalent.
2. Definitions and helpful results
We assume that the reader is familiar with the fundamentals of universal algebra as given in [3], [13] or [1]. Our notation generally follows that in [13], [1] and [9]. If ${\mathbf{A}}$ is an algebra, then $\operatorname{Con}({\mathbf{A}})$ denotes its congruence lattice. The smallest and largest congruences of ${\mathbf{A}}$ are the diagonal $0_{A}:=\{(a,a)\,:\,\text{$a\in A$}\}$ and the full congruence $1_{A}:=A^{2}$ respectively; they may be denoted 00 and $1$ if no confusion arises. If $\alpha,\beta\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\leq\beta$, then $\beta/\alpha$ denotes the congruence of ${\mathbf{A}}/\alpha$ corresponding to $\beta$ via the Correspondence Theorem ([13, Theorem 4.12] or [1, Theorem 3.6]). If in addition $\gamma,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\gamma\leq\delta$, then we write $(\alpha,\beta)\nearrow(\gamma,\delta)$ and say that $(\alpha,\beta)$ is perspective up to $(\gamma,\delta)$, or transposes up to $(\gamma,\delta)$, if $\beta\wedge\gamma=\alpha$ and $\beta\vee\gamma=\delta$. The notation $(\gamma,\delta)\searrow(\alpha,\beta)$ means the same thing. We write $\alpha\prec\beta$ and say that $\beta$ covers $\alpha$, or is an upper cover of $\alpha$, if $\alpha<\beta$ and there does not exist a congruence $\gamma$ satisfying $\alpha<\gamma<\beta$. A congruence $\alpha$ is minimal if it covers 00, and is completely meet-irreducible if $\alpha\neq 1$ and there exists $\alpha^{+}$ with $\alpha\prec\alpha^{+}$ and such that $\alpha<\beta\implies\alpha^{+}\leq\beta$ for all $\beta\in\operatorname{Con}({\mathbf{A}})$. When ${\mathbf{A}}$ is finite, we use the phrase “meet-irreducible” to mean the same thing. We say that ${\mathbf{A}}$ is subdirectly irreducible if 00 is completely meet-irreducible, in which case $0^{+}$ is called the monolith of ${\mathbf{A}}$. A subset $T\subseteq A$ is a transversal for a congruence $\alpha$ if it contains exactly one element from each $\alpha$-class.
If $f:A\rightarrow B$ is a function, then its graph is the set $\operatorname{graph}(f)=\{(a,f(a))\,:\,\text{$a\in A$}\}$. If $n>0$, then $[n]$ denotes $\{1,2,\ldots,n\}$. If $R\subseteq A_{1}\times\cdots\times A_{n}$ and $i,j\in[n]$, then $\operatorname{pr}_{i}(R)$ denotes the projection of $R$ onto its $i$-th coordinate, and $\operatorname{pr}_{i,j}(R)$ denotes $\{(a_{i},a_{j})\,:\,\text{$(a_{1},\ldots,a_{n})\in R$}\}$.
We follow [9] and refer to a set of operation symbols with assigned arities as a signature. Every algebra comes equipped with a signature, which indexes the basic operations of the algebra. Terms are formal recipes for constructing new operations from the basic operations via composition and variable manipulations; see [13] or [1] or any textbook on first-order logic. We will not distinguish between terms and the term operations they define in an algebra, except when the distinction is crucial. A polynomial of an algebra ${\mathbf{A}}$ is any operation on $A$ having the form $t(x_{1},\ldots,x_{n},\mathbf{c})$ where $t$ is an $(n+k)$-ary term in the signature of ${\mathbf{A}}$ and $\mathbf{c}\in A^{k}$. $\operatorname{Pol}_{n}({\mathbf{A}})$ is the set of all $n$-ary polynomials of ${\mathbf{A}}$.
If ${\mathbf{A}}$ is an algebra, a term $t(x_{1},\ldots,x_{n})$ of ${\mathbf{A}}$ is idempotent if ${\mathbf{A}}$ satisfies the identity $t(x,\ldots,x)\approx x$, and is a Taylor term if it is idempotent, $n>1$, and for each $i\in[n]$, ${\mathbf{A}}$ satisfies an identity of the form
$t(u_{1},\ldots,u_{n})\approx t(v_{1},\ldots,v_{n})$
where each $u_{j}$ and $v_{k}$ is the variable $x$ or $y$, and $\{u_{i},v_{i}\}=\{x,y\}$. An algebra is said to be Taylor if it has a Taylor term. A particularly important example of a Taylor term is a weak near-unanimity term (WNU), which is an $n$-ary idempotent term $w(x_{1},\ldots,x_{n})$ with $n>1$ which satisfies the identities
$w(y,x,x,\ldots,x)\approx w(x,y,x,\ldots,x)\approx w(x,x,y,\ldots,x)\approx\cdots\approx w(x,\ldots,x,y).$
Another important example of a Taylor term is a Maltsev term; this is a ternary term $p(x,y,z)$ satisfying the identities
(2.1) $p(x,x,y)\approx y\approx p(y,x,x).$
The identities (2.1) are called the Maltsev identities. Any ternary operation (whether a term or not) satisfying them is called a Maltsev operation.
A variety is a class of algebras (in a common signature) which is closed under subalgebras, homomorphic images, and direct products of arbitrary (including infinite) families of algebras. $\mathsf{HSP}({\mathbf{A}})$ denotes the smallest variety containing ${\mathbf{A}}$. A term is a Taylor term, or a WNU term, for a variety if it is such for every algebra in the variety. Because the definitions of Taylor terms and WNU terms are given in terms of satisfied identities, a Taylor or WNU term for an algebra ${\mathbf{A}}$ is automatically a Taylor or WNU term for the variety $\mathsf{HSP}({\mathbf{A}})$.
Before defining “weak difference term,” we recall the ternary centralizer relation on congruences and the notion of abelian congruences. Given a non-empty set $A$, let $A^{2\times 2}$ denote the set of all $2\times 2$ matrics over $A$. If ${\mathbf{A}}$ is an algebra, let ${\mathbf{A}}^{2\times 2}$ denote the algebra with universe $A^{2\times 2}$ which is isomorphic to ${\mathbf{A}}^{4}$ via the bijection
$\begin{pmatrix}a_{1}&a_{3}\\ a_{2}&a_{4}\end{pmatrix}\mapsto(a_{1},a_{2},a_{3},a_{4}).$
Definition 2.1.
Suppose ${\mathbf{A}}$ is an algebra and $\theta,\varphi\in\operatorname{Con}({\mathbf{A}})$. $M(\theta,\varphi)$ is the subuniverse of ${\mathbf{A}}^{2\times 2}$ generated by the set
$X(\theta,\varphi):=\left\{\begin{pmatrix}c&c\\ d&d\end{pmatrix}:(c,d)\in\theta\right\}\cup\left\{\begin{pmatrix}a&b\\ a&b\end{pmatrix}:(a,b)\in\varphi\right\}.$
The matrices in $M(\theta,\varphi)$ are called $(\theta,\varphi)$-matrices.
Definition 2.2.
Suppose $\theta,\varphi,\delta\in\operatorname{Con}({\mathbf{A}})$. We say that $\varphi$ centralizes $\theta$ modulo $\delta$, and write $C(\varphi,\theta;\delta)$, if any of the following equivalent conditions holds:
  1. (1)
    For every matrix in $M(\varphi,\theta)$, if one row is in $\delta$, then so is the other row.
  2. (2)
    For every matrix in $M(\theta,\varphi)$, if one column is in $\delta$, then so is the other column.
  3. (3)
    For every $(1+n)$-ary term $t(x,y_{1},\ldots,y_{n})$ and all $(a,b)\in\varphi$ and $(c_{j},d_{j})\in\theta$ ($j\in[n]$),
    if $t(a,\mathbf{c})\stackrel{{\scriptstyle\delta}}{{\equiv}}t(a,\mathbf{d})$, then $t(b,\mathbf{c})\stackrel{{\scriptstyle\delta}}{{\equiv}}t(b,\mathbf{d})$.
Let $\theta,\delta$ be congruences of an algebra ${\mathbf{A}}$. The centralizer (or annihilator) of $\theta$ modulo $\delta$, denoted $(\delta:\theta)$, is the unique largest congruence $\varphi$ for which $C(\varphi,\theta;\delta)$ holds. We say that $\theta$ is abelian if $C(\theta,\theta;0)$ holds; equivalently, if $\theta\leq(0:\theta)$. We say that ${\mathbf{A}}$ is abelian if $1_{A}$ is abelian. More generally, if $\theta,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\delta\leq\theta$, then we say that $\theta$ is abelian modulo $\delta$ if any of the following equivalent conditions holds: (i) $\theta/\delta$ is an abelian congruence of ${\mathbf{A}}/\delta$; (ii) $C(\theta,\theta;\delta)$ holds; (iii) $\theta\leq(\delta:\theta)$.
Definition 2.3.
Let $\mathcal{V}$ be a variety, ${\mathbf{A}}\in\mathcal{V}$, and $d(x,y,z)$ a ternary term in the signature of $\mathcal{V}$.
  1. (1)
    $d$ is a weak difference term for ${\mathbf{A}}$ if $d$ is idempotent and for every pair $\delta,\theta$ of congruences with $\delta\leq\theta$ and $\theta/\delta$ abelian, we have
    (2.2) $d(a,a,b)\stackrel{{\scriptstyle\delta}}{{\equiv}}b\stackrel{{\scriptstyle\delta}}{{\equiv}}d(b,a,a)$ for all $(a,b)\in\theta$.
  2. (2)
    $d$ is a weak difference term for $\mathcal{V}$ if it is a weak difference term for every algebra in $\mathcal{V}$.
Note in particular that if $d$ is a weak difference term for ${\mathbf{A}}$ and $\theta$ is an abelian congruence, then setting $\delta=0$ in (2.2) gives that the restriction of $d$ to any $\theta$-class is a Maltsev operation on that class. In fact, $d$ induces an abelian group operation on each $\theta$-class in this case, by the following result of Gumm [5] and Herrmann [7].
Definition 2.4.
Suppose ${\mathbf{A}}$ is an algebra having a weak difference term $d(x,y,z)$, and $\theta$ is an abelian congruence of ${\mathbf{A}}$. Given $e\in A$, let $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ denote the algebra $(e/\theta,+,e)$ whose universe is the $\theta$-class containing $e$ and whose two operations are the binary operation $x+y:=d(x,e,y)$ and the constant $e$.
Lemma 2.5 (essentially [5, 7]; cf. [16, Lemma 3.5]).
Suppose ${\mathbf{A}},d,\theta$ are as in the previous definition and $e\in A$. Then $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ is an abelian group with zero element $e$. Moreover, we have $-x=d(e,x,e)$ and $d(x,y,z)=x-y+z$ for all $x,y,z\in e/\theta$.
The next result is folklore.
Lemma 2.6 (cf. [16, Lemma 3.1]).
Suppose ${\mathbf{A}}$ is an algebra and $\rho$ is a reflexive subuniverse of ${\mathbf{A}}^{2}$. Suppose ${\mathbf{A}}$ has a ternary term $d(x,y,z)$ such that for all $(a,b)\in\rho$ we have $d(a,a,b)=b$ and $d(a,b,b)=a$. Then $\rho\in\operatorname{Con}({\mathbf{A}})$.
The next lemma will be needed in Section 6.
Lemma 2.7.
Suppose ${\mathbf{A}}$ is an algebra having a weak difference term, $\mu\in\operatorname{Con}({\mathbf{A}})$ is an abelian minimal congruence, and $S\leq{\mathbf{A}}$ is a subuniverse of ${\mathbf{A}}$ which is a transversal for $\mu$. Then $S$ is a maximal proper subuniverse of ${\mathbf{A}}$.
Proof.
Clearly $S\neq A$ as $\mu\neq 0_{A}$. Let $\pi:{\mathbf{A}}\rightarrow{\mathbf{S}}$ be the homomorphism given by letting $\pi(x)$ be the unique element $y\in S$ satisfying $(x,y)\in\mu$. Fix $a\in A\setminus S$. Let $\rho=\operatorname{Sg}^{{\mathbf{A}}^{2}}(\{(a,\pi(a))\}\cup 0_{A})$. $\rho$ is a reflexive subuniverse of ${\mathbf{A}}^{2}$ satisfying $0_{A}\neq\rho\subseteq\mu$. Thus by Lemma 2.6 and minimality of $\mu$, we get $\rho=\mu$. Now let $b\in A$ be arbitrary. Then $(b,\pi(b))\in\mu=\rho$, so we can pick a term $t(x,\mathbf{y})$ and a tuple $\mathbf{c}=(c_{1},\ldots,c_{n})$ of elements of $A$ such that
$(b,\pi(b))=t^{{\mathbf{A}}^{2}}((a,\pi(a)),(c_{1},c_{1}),\ldots,(c_{n},c_{n}))=(t(a,\mathbf{c}),t(\pi(a),\mathbf{c})).$
Let $\mathbf{u}$ be the tuple in $S$ obtained by applying $\pi$ coordinatewise to $\mathbf{c}$. Observe that $\pi(b)=\pi(t(a,\mathbf{c}))=t(\pi(a),\mathbf{u})$, so
$t(\underline{\pi(a)},\mathbf{c})=t(\underline{\pi(a)},\mathbf{u}).$
As $\mathbf{c},\mathbf{u}$ are coordinatewise in $\mu$, $(a,\pi(a))\in\mu$, and $\mu$ is abelian, we can use Definition 2.2(3) and replace the underlined instances of $\pi(a)$ with $a$ to get $t(a,\mathbf{c})=t(a,\mathbf{u})$, proving $b=t(a,\mathbf{u})\in\operatorname{Sg}(S\cup\{a\})$. As $b$ was arbitrary, we get $\operatorname{Sg}(S\cup\{a\})=A$. ∎
Some proofs in Sections 5 and 6 use tame congruence theory, of which we will give only a cursory overview. Given a finite algebra ${\mathbf{A}}$, to each pair $(\alpha,\beta)$ of congruences of ${\mathbf{A}}$ with $\alpha\prec\beta$, the theory assigns one of five “types” from $\{\mathbf{1},\mathbf{2},\mathbf{3},\mathbf{4},\mathbf{5}\}$. The theory also defines “$(\alpha,\beta)$-minimal sets,” which are special subsets of the universe $A$, and “$(\alpha,\beta)$-traces,” which are sets of the form $U\cap C$ where $U$ is an $(\alpha,\beta)$-minimal set, $C$ is a $\beta$-class, and $(U\cap C)^{2}\nsubseteq\alpha$. Most of what we will need is contained in the following two results.
Proposition 2.8 ([8]).
Suppose ${\mathbf{A}}$ is a finite algebra and $\alpha,\beta\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\prec\beta$.
  1. (1)
    There exists an $(\alpha,\beta)$-minimal set.
  2. (2)
    $\beta/\alpha$ is nonabelian if and only if $\operatorname{typ}(\alpha,\beta)\in\{\mathbf{3},\mathbf{4},\mathbf{5}\}$.
  3. (3)
    Suppose $\beta/\alpha$ is nonabelian and $U$ is an $(\alpha,\beta)$-minimal set. Then there exists a unique $\beta$-class $C$ such that $U\cap C$ is an $(\alpha,\beta)$-trace. Moreover, letting $N:=U\cap C$, there exist $1\in N$, a unary polynomial $e(x)\in\operatorname{Pol}_{1}({\mathbf{B}})$, and a binary polynomial $p(x,y)\in\operatorname{Pol}_{2}({\mathbf{A}})$ satisfying:
    1. (a)
      $(N\setminus\{1\})^{2}\subseteq\alpha$.
    2. (b)
      $e(A)=U$ and $e(x)=x$ for all $x\in U$.
    3. (c)
      $p(x,1)=p(1,x)=p(x,x)=x$ for all $x\in U$.
    4. (d)
      $p(x,o)\stackrel{{\scriptstyle\alpha}}{{\equiv}}p(o,x)\stackrel{{\scriptstyle\alpha}}{{\equiv}}x$ for all $x\in U\setminus\{1\}$ and all $o\in N\setminus\{1\}$.
Proof.
(1) follows from [8, Theorem 5.7(1) and Theorem 2.8(2)]. (2) follows from [8, Theorem 5.7(1,2)]. (2) can be deduced from [8, Theorem 5.7(1) and Lemmas 2.13(3), 4.15 and 4.17]. ∎
A variety is locally finite if its finitely generated algebras are always finite. As an example, $\mathsf{HSP}({\mathbf{A}})$ is locally finite whenever ${\mathbf{A}}$ is finite. We say that a variety omits type $\mathbf{i}$ if no finite algebra in the variety has a pair of congruences $\alpha\prec\beta$ with $\operatorname{typ}(\alpha,\beta)=\mathbf{i}$.
Theorem 2.9.
For a locally finite variety $\mathcal{V}$, the following are equivalent:
  1. (1)
    $\mathcal{V}$ has a Taylor term.
  2. (2)
    $\mathcal{V}$ has a WNU term.
  3. (3)
    $\mathcal{V}$ has a weak difference term.
  4. (4)
    $\mathcal{V}$ omits type $\mathbf{1}$.
Proof.
This follows by combining [8, Theorem 9.6], [14, Corollary 5.3], [12, Theorem 2.2], and [10, Theorem 4.8]. ∎
3. Similarity in varieties with a weak difference term
In this section we list the definitions and results about similarity from [16] which we will use in Section 6.
Definition 3.1.
Suppose ${\mathbf{A}}$ is an algebra and $\theta,\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\theta\leq\alpha$.
  1. (1)
    ${\mathbf{A}}(\theta)$ denotes $\theta$ viewed as a subalgebra of ${\mathbf{A}}^{2}$.
  2. (2)
    $\Delta_{\theta,\alpha}$ denotes the congruence of ${\mathbf{A}}(\theta)$ generated by $\{((a,a),(b,b))\,:\,\text{$(a,b)\in\alpha$}\}$.
Lemma 3.2.
Let ${\mathbf{A}}$ be an algebra and $\delta,\theta,\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\theta\leq\alpha$ and $C(\alpha,\theta;\delta)$. Then for all $((a,a^{\prime}),(b,b^{\prime}))\in\Delta_{\theta,\alpha}$, $(a,a^{\prime})\in\delta\iff(b,b^{\prime})\in\delta$.
Proof.
$\Delta_{\theta,\alpha}$ is the transitive closure of $\{((a,a^{\prime}),(b,b^{\prime}))\,:\,\text{${\scriptsize{\begin{pmatrix}a\phantom{{}^{\prime}}&b\phantom{{}^{\prime}}\\ a^{\prime}&b^{\prime}\end{pmatrix}}}\in M(\theta,\alpha)$}\}$ (see e.g. [16, Lemma 4.4(1)]). Hence the result follows by $C(\alpha,\theta;\delta)$ and Definition 2.2(2). ∎
We are mainly interested in $\Delta_{\theta,\alpha}$ when $\theta$ is abelian and $\alpha=(0:\theta)$. In this context we use the following notation.
Definition 3.3.
Suppose ${\mathbf{A}}$ is an algebra, $\theta\in\operatorname{Con}({\mathbf{A}})$ is abelian, and $\alpha=(0:\theta)$.
  1. (1)
    $\overline{\alpha}$ denotes the set $\{((a,a^{\prime}),(b,b^{\prime}))\in A(\theta)^{2}\,:\,\text{$(a,b)\in\alpha$}\}$, which is a congruence of ${\mathbf{A}}(\theta)$ satisfying $\Delta_{\theta,\alpha}\leq\overline{\alpha}$.
  2. (2)
    $D({\mathbf{A}},\theta)$ denotes the quotient algebra ${\mathbf{A}}(\theta)/\Delta_{\theta,\alpha}$.
Theorem 3.4 ([16, Corollary 4.8]).
Suppose ${\mathbf{A}}$ belongs to a variety with a weak difference term and $\theta$ is an abelian minimal congruence of ${\mathbf{A}}$. Let $\alpha=(0:\theta)$ and $\varphi=\overline{\alpha}/\Delta_{\theta,\alpha}$. Then $D({\mathbf{A}},\theta)$ is subdirectly irreducible with abelian monolith $\varphi$. Moreover, $(0:\varphi)=\varphi$, and there exist a subuniverse $D^{o}\leq D({\mathbf{A}},\theta)$, a surjective homomorphism $h:{\mathbf{A}}(\theta)\rightarrow D({\mathbf{A}},\theta)$, and an isomorphism $h^{\ast}:{\mathbf{A}}/\alpha\cong D({\mathbf{A}},\theta)/\varphi$ such that:
  1. (1)
    $D^{o}$ is a transversal for $\varphi$.
  2. (2)
    $h^{-1}(D^{o})=0_{A}$.
  3. (3)
    $h$ and $h^{\ast}$ are compatible in the following sense: for all $(a,b)\in\theta$, $h(a,b)/\varphi=h^{\ast}(a/\alpha)$.
Definition 3.5.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}}\in\mathcal{V}$ is subdirectly irreducible with monolith $\mu$. The algebra $D({\mathbf{A}})$ is defined as follows:
  1. (1)
    If $\mu$ is nonabelian, then $D({\mathbf{A}})={\mathbf{A}}$.
  2. (2)
    If $\mu$ is abelian, then $D({\mathbf{A}})=D({\mathbf{A}},\mu)$ as defined in Definition 3.3.
Definition 3.6.
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible. We say that ${\mathbf{A}}$ and ${\mathbf{B}}$ are similar, and write ${\mathbf{A}}\sim{\mathbf{B}}$, if $D({\mathbf{A}})\cong D({\mathbf{B}})$.
The following definition from [16] was motivated by Zhuk’s bridges [18].
Definition 3.7.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are subdirectly irreducible algebras in a common signature with monoliths $\mu,\kappa$ respectively. A proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$ is a subuniverse $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ satisfying
  1. (B1)
    $\operatorname{pr}_{1,2}(T)=\mu$ and $\operatorname{pr}_{3,4}(T)=\kappa$.
  2. (B2)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $a_{1}=a_{2}$ if and only if $b_{1}=b_{2}$.
  3. (B3)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $(a_{i},a_{i},b_{i},b_{i})\in T$ for $i=1,2$.
Theorem 3.8 ([16, Theorem 7.8]).
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible. The following are equivalent:
  1. (1)
    ${\mathbf{A}}\sim{\mathbf{B}}$.
  2. (2)
    There exist an algebra ${\mathbf{C}}\in\mathcal{V}$, surjective homomorphisms $f_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $f_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$, and congruences $\psi,\tau\in\operatorname{Con}({\mathbf{C}})$ with $\psi<\tau$, such that, letting $\delta_{i}=\ker(f_{i})$ and letting $\delta^{+}_{i}$ denote the unique upper cover of $\delta_{i}$ in $\operatorname{Con}({\mathbf{C}})$ for $i=1,2$, we have $(\delta_{1},\delta^{+}_{1})\searrow(\psi,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$.
  3. (3)
    There exists a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$.
Corollary 3.9 ([16, Corollary 7.9]).
Suppose $\mathcal{V}$ is a variety with a weak difference term, and ${\mathbf{A}}\in\mathcal{V}$ is subdirectly irreducible with abelian monolith $\mu$. Setting $\alpha=(0:\mu)$ and $\Delta=\Delta_{\mu,\alpha}$, the set
$T_{{\mathbf{A}}}^{D}:=\{(a,b,(a,e)/\Delta,(b,e)/\Delta)\,:\,\text{$a\stackrel{{\scriptstyle\mu}}{{\equiv}}b\stackrel{{\scriptstyle\mu}}{{\equiv}}e$}\}$
is a proper bridge from ${\mathbf{A}}$ to $D({\mathbf{A}})$.
4. Zhuk’s bridges
In his solution to the Constraint Satisfaction Problem Dichotomy Conjecture [18], D. Zhuk defined and used to great effect certain devices which he called “bridges.” In this section we present Zhuk’s bridges, in slightly greater generality than Zhuk’s original setting. To help clarify their development, we introduce some new terminology and recast some of Zhuk’s basic proofs.
Definition 4.1.
Let ${\mathbf{A}}$ be an algebra and $\rho\in\operatorname{Con}({\mathbf{A}})$. A subuniverse $R\leq{\mathbf{A}}^{2}$ is said to be stable under $\rho$ [18], or $\rho$-saturated [11], or $\rho$-closed [8], if $R=\rho\circ R\circ\rho$.
Definition 4.2.
An anchor is a triple $({\mathbf{A}},\rho,R)$ where ${\mathbf{A}}$ is a finite algebra, $\rho\in\operatorname{Con}({\mathbf{A}})$, $R\leq{\mathbf{A}}^{2}$, $R$ is stable under $\rho$, and $R$ properly contains $\rho$.
Definition 4.3.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are finite algebras in a common signature and $({\mathbf{A}},\rho,R)$, $({\mathbf{B}},\sigma,S)$ are anchors. A bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{B}},\sigma,S)$ is a subuniverse $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ satisfying
  1. (B0)
    $T$ is “stable under $\rho$ in its first two coordinates and stable under $\sigma$ in its last two coordinates.” That is, for all $(a_{1},a_{2},b_{1},b_{2}),(a_{1}^{\prime},a_{2}^{\prime},b_{1}^{\prime},b_{2}^{\prime})\in A\times A\times B\times B$ such that $(a_{i},a_{i}^{\prime})\in\rho$ and $(b_{i},b_{i}^{\prime})\in\sigma$ for $i=1,2$, we have
    $(a_{1},a_{2},b_{1},b_{2})\in T\iff(a_{1}^{\prime},a_{2}^{\prime},b_{1}^{\prime},b_{2}^{\prime})\in T.$
  2. (B1)
    $\operatorname{pr}_{1,2}(T)=R$ and $\operatorname{pr}_{3,4}(T)=S$.
  3. (B2)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $(a_{1},a_{2})\in\rho\iff(b_{1},b_{2})\in\sigma$.
Condition (B0), together with the requirement of stability in the definition of anchors, allow for “faithfully modding out by $\rho$ and $\sigma$.” That is, if $({\mathbf{A}},\rho,R)$ and $({\mathbf{B}},\sigma,S)$ are general anchors and $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ with $T,\rho,\sigma$ satisfying (B0), then setting $\overline{{\mathbf{A}}}:={\mathbf{A}}/\rho$ and $\overline{{\mathbf{B}}}:={\mathbf{B}}/\sigma$, we can form quotients $\overline{R}\leq(\overline{{\mathbf{A}}})^{2}$, $\overline{S}\leq(\overline{{\mathbf{B}}})^{2}$, and $\overline{T}\leq\overline{{\mathbf{A}}}\times\overline{{\mathbf{A}}}\times\overline{{\mathbf{B}}}\times\overline{{\mathbf{B}}}$ in the obvious way, namely,
$\overline{R}$ =$=\{(a/\rho,\,a^{\prime}/\rho)\,:\,\text{$(a,a^{\prime})\in R$}\}$
$\overline{S}$ =$=\{(b/\sigma,\,b^{\prime}/\sigma)\,:\,\text{$(b,b^{\prime})\in S$}\}$
$\overline{T}$ =$=\{(a/\rho,\,a^{\prime}/\rho,\,b/\sigma,\,b^{\prime}/\sigma)\,:\,\text{$(a,a^{\prime},b,b^{\prime})\in T$}\}.$
Then $(\overline{{\mathbf{A}}},0_{\overline{A}},\overline{R})$ and $(\overline{{\mathbf{B}}},0_{\overline{B}},\overline{S})$ are anchors, and $T$ is a bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{B}},\sigma,S)$ if and only if $\overline{T}$ is a bridge from $(\overline{{\mathbf{A}}},0_{\overline{A}},\overline{R})$ to $(\overline{{\mathbf{B}}},0_{\overline{B}},\overline{S})$. Moreover, $R$, $S$ and $T$ are recoverable from $\overline{R}$, $\overline{S}$ and $\overline{T}$ respectively, as the pre-images under the natural maps ${\mathbf{A}}^{2}\rightarrow(\overline{{\mathbf{A}}})^{2}$, ${\mathbf{B}}^{2}\rightarrow(\overline{{\mathbf{B}}})^{2}$, and ${\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}\rightarrow\overline{{\mathbf{A}}}\times\overline{{\mathbf{A}}}\times\overline{{\mathbf{B}}}\times\overline{{\mathbf{B}}}$ of $\overline{R}$, $\overline{S}$ and $\overline{T}$ respectively.
As the next definition suggests, we can view bridges as the arrows in a category whose objects are the anchors. Moreover, every bridge has a natural “converse.”
Definition 4.4.
Suppose ${\mathbf{A}},{\mathbf{B}},{\mathbf{C}}$ are finite algebras in a common signature and $({\mathbf{A}},\rho,R)$, $({\mathbf{B}},\sigma,S),({\mathbf{C}},\upsilon,U)$ are anchors. Let $T$ be a bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{B}},\sigma,S)$ and let $T^{\prime}$ be a bridge from $({\mathbf{B}},\sigma,S)$ to $({\mathbf{C}},\upsilon,U)$.
  1. (1)
    The identity bridge from $({\mathbf{A}},\rho,R)$ to itself is the relation given by
    $I_{({\mathbf{A}},\rho,R)}=\{(a_{1},a_{2},b_{1},b_{2})\in A^{4}\,:\,\text{$(a_{1},a_{2}),(b_{1},b_{2})\in R$ and $(a_{1},b_{1}),(a_{2},b_{2})\in\rho$}\}.$
  2. (2)
    The converse of $T$ is the subuniverse $T^{\cup}\leq{\mathbf{B}}\times{\mathbf{B}}\times{\mathbf{A}}\times{\mathbf{A}}$ given by
    $T^{\cup}=\{(b_{1},b_{2},a_{1},a_{2})\,:\,\text{$(a_{1},a_{2},b_{1},b_{2})\in T$}\}.$
  3. (3)
    The composition $T\circ T^{\prime}$ is the subuniverse $T\circ T^{\prime}\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{C}}\times{\mathbf{C}}$ given by
    $T\circ T^{\prime}=\{(a_{1},a_{2},c_{1},c_{2})\,:\,\text{$\exists b_{1},b_{2}\in B$ with $(a_{1},a_{2},b_{1},b_{2})\in T$ and $(b_{1},b_{2},c_{1},c_{2})\in T^{\prime}$}\}.$
It is easy to check that identity bridges are bridges, the converse of a bridge is a bridge, and the composition of bridges is a bridge.
An important invariant of a bridge is its “trace.”
Definition 4.5.
Suppose $T$ is a bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{B}},\sigma,S)$.
  1. (1)
    The trace of $T$, denoted $\operatorname{tr}(T)$, is the subuniverse of ${\mathbf{A}}\times{\mathbf{B}}$ defined by $\operatorname{tr}(T)=\{(a,b)\,:\,\text{$(a,a,b,b)\in T$}\}$.
  2. (2)
    When ${\mathbf{B}}={\mathbf{A}}$, we say that $T$ is reflexive if $0_{A}\subseteq\operatorname{tr}(T)$.
Zhuk [18] denoted $\operatorname{tr}(T)$ by $\widetilde{T}$ (and did not call it a “trace”). One can check that $\operatorname{tr}(I_{({\mathbf{A}},\rho,R)})=\rho$, $\operatorname{tr}(T^{\cup})=\operatorname{tr}(T)^{-1}$, and $\operatorname{tr}(T\circ T^{\prime})=\operatorname{tr}(T)\circ\operatorname{tr}(T^{\prime})$.
Remark 4.6.
In his original definition, Zhuk viewed a bridge $T$ as being “from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$,” not “from $({\mathbf{A}},\rho,R)$ to $({\mathbf{B}},\sigma,S)$.” The relations $R$ and $S$ were left implicit, though recoverable from $T$ via the equations $R=\operatorname{pr}_{1,2}(T)$ and $S=\operatorname{pr}_{3,4}(T)$.
Definition 4.7.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are finite algebras in a common signature, $\rho\in\operatorname{Con}({\mathbf{A}})$ with $\rho\neq 1_{A}$, and $\sigma\in\operatorname{Con}({\mathbf{B}})$ with $\sigma\neq 1_{B}$. A (Zhuk) bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$ is any bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{B}},\sigma,S)$, for some subuniverses $R\leq{\mathbf{A}}^{2}$ and $S\leq{\mathbf{B}}^{2}$ for which $({\mathbf{A}},\rho,R)$ and $({\mathbf{B}},\sigma,S)$ are anchors.
Zhuk’s bridges cannot be viewed as arrows in a category whose objects are pairs $({\mathbf{A}},\rho)$, as composition is then not always defined. Happily, Zhuk only ever used his bridges between pairs $({\mathbf{A}},\rho),({\mathbf{B}},\sigma)$ for which there are canonical choices for the relations $R$ and $S$.
Definition 4.8 (Zhuk [18]).
Let ${\mathbf{A}}$ be a finite algebra and $\rho\in\operatorname{Con}({\mathbf{A}})$.
  1. (1)
    $\rho$ is irreducible if $\rho\neq 1_{A}$ and $\rho$ cannot be written as the intersection of two $\rho$-saturated subuniverses of ${\mathbf{A}}^{2}$ both properly containing $\rho$.
  2. (2)
    If $\rho$ is irreducible, then $\rho^{\ast}$ denotes the unique smallest $\rho$-saturated subuniverse of ${\mathbf{A}}^{2}$ properly containing $\rho$.
The following fact is easily proved.
Lemma 4.9.
If ${\mathbf{A}}$ is finite and $\rho\in\operatorname{Con}({\mathbf{A}})$ is irreducible, then $\rho$ is meet-irreducible in $\operatorname{Con}({\mathbf{A}})$; its unique upper cover $\rho^{+}$ in $\operatorname{Con}({\mathbf{A}})$ is the transitive closure of $\rho^{\ast}$.
The next lemma explains the tight connection between Zhuk’s bridges and bridges as we defined them in Definition 4.3, when the congruences are irreducible.
Lemma 4.10 (Zhuk [18]).
Suppose ${\mathbf{A}},{\mathbf{B}}$ are finite algebras in a common signature, $\rho\in\operatorname{Con}({\mathbf{A}})$, and $\sigma\in\operatorname{Con}({\mathbf{B}})$. If $\rho$ and $\sigma$ are irreducible, then every bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$ contains a bridge from $({\mathbf{A}},\rho,\rho^{\ast})$ to $({\mathbf{B}},\sigma,\sigma^{\ast})$ with the same trace.
Proof (Zhuk [18]).
Suppose $T$ is a bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{B}},\sigma,S)$. Define $T_{1}=\{(a_{1},a_{2},b_{1},b_{2})\in T\,:\,\text{$(a_{1},a_{2})\in\rho^{\ast}$}\}$. Let $S_{1}=\operatorname{pr}_{3,4}(T_{1})$. As $\rho^{\ast}\subseteq R=\operatorname{pr}_{1,2}(T)$, we get $\rho^{\ast}=\operatorname{pr}_{1,2}(T_{1})$. In particular, there exists $(a_{1},a_{2},b_{1},b_{2})\in T_{1}$ with $(a_{1},a_{2})\not\in\rho$. By property (B2), this implies $(b_{1},b_{2})\not\in\sigma$. Hence $S_{1}\neq\sigma$. On the other hand, $\sigma\subseteq S=\operatorname{pr}_{3,4}(T)$ and (B2) imply $\sigma\subseteq S_{1}$, and (B0) implies $S_{1}$ is stable under $\sigma$. As $\sigma$ is irreducible, these facts imply $\sigma^{\ast}\subseteq S_{1}$. It can be checked that $T_{1}$ is a bridge from $({\mathbf{A}},\rho,\rho^{\ast})$ to $({\mathbf{B}},\sigma,S_{1})$ with $\operatorname{tr}(T_{1})=\operatorname{tr}(T)$.
Next let $T_{2}=\{(a_{1},a_{2},b_{1},b_{2})\in T_{1}\,:\,\text{$(b_{1},b_{2})\in\sigma^{\ast}$}\}$. A similar argument shows $T_{2}$ is a bridge from $({\mathbf{A}},\rho,R_{2})$ to $({\mathbf{B}},\sigma,\sigma^{\ast})$ with $\operatorname{tr}(T_{2})=\operatorname{tr}(T_{1})$, where $R_{2}=\operatorname{pr}_{1,2}(T_{2})$. But clearly $\rho^{\ast}\subseteq R_{2}\subseteq\operatorname{pr}_{1,2}(T_{1})=\rho^{\ast}$, i.e., $R_{2}=\rho^{\ast}$ as required. ∎
Definition 4.11.
Given an anchor $({\mathbf{A}},\rho,R)$, we let $\operatorname{ref.tr}(\rho,R)$ denote the set
$\operatorname{ref.tr}(\rho,R)=\{\operatorname{tr}(T)\,:\,\text{$T$ is a reflexive bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{A}},\rho,R)$}\}.$
Note that each member of $\operatorname{ref.tr}(\rho,R)$ is a subuniverse of ${\mathbf{A}}^{2}$ containing $\rho$, by (B0). In addition, $\operatorname{ref.tr}(\rho,R)$ is nonempty and closed under inversion and composition by the comments following Definitions 4.4 and 4.5. It follows from this and finiteness that $\operatorname{ref.tr}(\rho,R)$ contains a unique maximal member; and this unique maximal member is a congruence containing $\rho$. The following definition and lemma record this observation.
Definition 4.12.
If $({\mathbf{A}},\rho,R)$ is an anchor, then $\operatorname{Opt}(\rho,R)$ denotes the unique maximal member of $\operatorname{ref.tr}(\rho,R)$.
Lemma 4.13.
For any anchor $({\mathbf{A}},\rho,R)$ we have $\rho\leq\operatorname{Opt}(\rho,R)\in\operatorname{Con}({\mathbf{A}})$.
Definition 4.14 (Zhuk [18]).
Suppose ${\mathbf{A}}$ is a finite algebra and $\rho\in\operatorname{Con}({\mathbf{A}})$ is irreducible.
  1. (1)
    $\operatorname{Opt}(\rho)$ denotes $\operatorname{Opt}(\rho,\rho^{\ast})$.
  2. (2)
    A bridge $T$ from $({\mathbf{A}},\rho)$ to $({\mathbf{A}},\rho)$ is optimal if $\operatorname{tr}(T)=\operatorname{Opt}(\rho)$.
Zhuk defined $\operatorname{Opt}(\rho)$ in a slightly different way: he let $\operatorname{Opt}(\rho)$ be the unique maximal member of the set
$\{\operatorname{tr}(T)\,:\,\text{$T$ is a reflexive bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{A}},\rho)$}\}.$
In fact, this set is identical to $\operatorname{ref.tr}(\rho,\rho^{\ast})$ by Lemma 4.10, so Zhuk’s and our definitions of $\operatorname{Opt}(\rho)$ are equivalent.
The following fact can be extracted from Zhuk’s proof of [18, Corollary 7.24.1].
Lemma 4.15.
Suppose $({\mathbf{A}},\rho,R)$ and $({\mathbf{B}},\sigma,S)$ are anchors and $T$ is a bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{B}},\sigma,S)$. Let $\alpha=\operatorname{Opt}(\rho,R)$ and $\beta=\operatorname{Opt}(\sigma,S)$. Then $\operatorname{tr}(T)$ induces an isomorphism $\gamma:{{\mathbf{A}}/\alpha}\cong{\mathbf{B}}/\beta$ defined by
$\gamma(a/\alpha)=b/\beta\iff(a,b)\in\alpha\circ\operatorname{tr}(T)\circ\beta.$
Proof.
Suppose $(a,b),(a^{\prime},b^{\prime})\in\operatorname{tr}(T)$. It suffices to show $(a,a^{\prime})\in\alpha\iff(b,b^{\prime})\in\beta$. Assume $(b,b^{\prime})\in\beta$. Let $T_{0}$ be a reflexive bridge from $({\mathbf{B}},\sigma,S)$ to itself satisfying $\operatorname{tr}(T_{0})=\beta$. Let $T_{1}=T\circ T_{0}$ and $T_{2}=T_{1}\circ T_{1}^{\cup}$. Then $(a,b^{\prime}),(a^{\prime},b^{\prime})\in\operatorname{tr}(T)\circ\beta=\operatorname{tr}(T_{1})$ and thus $(a,a^{\prime})\in\operatorname{tr}(T_{1})\circ\operatorname{tr}(T_{1})^{-1}=\operatorname{tr}(T_{2})$. Since $T_{2}$ is a reflexive bridge from $({\mathbf{A}},\rho,R)$ to itself, we get $\operatorname{tr}(T_{2})\subseteq\operatorname{Opt}(\rho,R)=\alpha$, so $(a,a^{\prime})\in\alpha$. Thus we have proved $(b,b^{\prime})\in\beta$ implies $(a,a^{\prime})\in\alpha$. A similar proof shows the opposite implication. ∎
Lemma 4.16.
Suppose $({\mathbf{A}},\rho,R)$ and $({\mathbf{A}},\sigma,S)$ are anchors on the same algebra. If there exists a reflexive bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{A}},\sigma,S)$, then $\operatorname{Opt}(\rho,R)=\operatorname{Opt}(\sigma,S)$.
Proof.
Let $T_{1}$ be a reflexive bridge from $({\mathbf{A}},\rho,R)$ to $({\mathbf{A}},\sigma,S)$. Let $T_{2}$ be a reflexive bridge from $({\mathbf{A}},\rho,R)$ to itself satisfying $\operatorname{tr}(T_{2})=\operatorname{Opt}(\rho,R)$. Let $T_{3}=T_{1}^{\cup}\circ T_{2}\circ T_{1}$. We have
$\operatorname{Opt}(\rho,R)$ $=0_{A}\circ\operatorname{Opt}(\rho,R)\circ 0_{A}$
$\subseteq\operatorname{tr}(T_{1}^{\cup})\circ\operatorname{tr}(T_{2})\circ\operatorname{tr}(T_{1})$
$=\operatorname{tr}(T_{3})$
$\subseteq\operatorname{Opt}(\sigma,S)$
where the last inclusion is because $T_{3}$ is a reflexive bridge from $({\mathbf{A}},\sigma,S)$ to itself. A symmetric argument shows $\operatorname{Opt}(\sigma,S)\subseteq\operatorname{Opt}(\rho,R)$. ∎
Definition 4.17 (cf. Zhuk [18]).
Suppose ${\mathbf{A}}$ is a finite algebra and $\rho,\sigma\in\operatorname{Con}({\mathbf{A}})$ are irreducible. We say that $\rho$ and $\sigma$ are adjacent if there exists a reflexive bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{A}},\sigma)$.
Corollary 4.18 (Zhuk [18, Lemma 6.4]).
Suppose ${\mathbf{A}}$ is a finite algebra and $\rho,\sigma\in\operatorname{Con}({\mathbf{A}})$ are irreducible and adjacent. Then $\operatorname{Opt}(\rho)=\operatorname{Opt}(\sigma)$.
Proof.
Let $T$ be a reflexive bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{A}},\sigma)$. By Lemma 4.10, we may assume that $T$ is from $({\mathbf{A}},\rho,\rho^{\ast})$ to $({\mathbf{A}},\sigma,\sigma^{\ast})$. Now apply Lemma 4.16. ∎
This completes our development of the basic terminology and results concerning bridges from [18].
5. Rooted bridges in locally finite Taylor varieties
In this short section we use tame congruence theory, first to characterize irreducible congruences in finite Taylor algebras, and secondly to extend the material from the previous section to (certain) bridges, which we call “rooted bridges,” between arbitrary meet-irreducible congruences of finite Taylor algebras.
Definition 5.1.
Suppose ${\mathbf{A}}$ is a finite algebra, $\rho\in\operatorname{Con}({\mathbf{A}})$ is meet-irreducible, and $\rho^{+}$ is its unique upper cover. $\operatorname{Cov}(\rho)$ denotes the set of minimal (under inclusion) $\rho$-saturated subuniverses $\tau$ of ${\mathbf{A}}^{2}$ satisfying $\rho\subset\tau\subseteq\rho^{+}$.
If $\rho$ is irreducible, then clearly $\operatorname{Cov}(\rho)=\{\rho^{\ast}\}$. However, it can happen that $\rho$ is meet-irreducible and $|\operatorname{Cov}(\rho)|=1$, yet $\rho$ is not irreducible; see Lemma 5.5.
Definition 5.2.
Suppose ${\mathbf{A}}$ and ${\mathbf{B}}$ are finite algebras in a common signature, $\rho\in\operatorname{Con}({\mathbf{A}})$, $\sigma\in\operatorname{Con}({\mathbf{B}})$, and $\rho$ and $\sigma$ are meet-irreducible. A bridge $T$ from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$ is rooted if there exist $\tau\in\operatorname{Cov}(\rho)$ and $\tau^{\prime}\in\operatorname{Cov}(\sigma)$ such that $T$ contains a bridge $T^{\prime}$ from $({\mathbf{A}},\rho,\tau)$ to $({\mathbf{B}},\sigma,\tau^{\prime})$ with $\operatorname{tr}(T^{\prime})=\operatorname{tr}(T)$.
Equivalently, a bridge $T$ from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$ is rooted if and only if the set
$T_{0}:=T\cap\{(a,a^{\prime},b,b^{\prime})\,:\,\text{$(a,a^{\prime})\in\rho^{+}$ and $(b,b^{\prime})\in\sigma^{+}$}\}$
satisfies $\operatorname{pr}_{1,2}(T_{0})\neq\rho$, where $\rho^{+},\sigma^{+}$ are the unique upper covers of $\rho,\sigma$ respectively.
Observe that if $\rho\in\operatorname{Con}({\mathbf{A}})$ and $\sigma\in\operatorname{Con}({\mathbf{B}})$ are irreducible, then every bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$ is rooted by Lemma 4.10. In extending Zhuk’s theory to meet-irreducible congruences, we will only ever consider rooted bridges.
Next, we introduce some notation and record two facts about $\operatorname{Cov}(\rho)$ given by tame congruence theory.
Definition 5.3.
Suppose ${\mathbf{A}}$ is a finite algebra, $\rho\in\operatorname{Con}({\mathbf{A}})$ is meet-irreducible, and $\rho^{+}$ is its unique upper cover. Let $\overline{\rho}$ denote the following subset of $\rho^{+}$:
$\overline{\rho}=\rho\circ\left(0_{A}\cup\bigcup\{N^{2}\,:\,\text{$N$ is a $(\rho,\rho^{+})$-trace}\}\right)\circ\rho.$
Proposition 5.4 ([8, Lemma 5.24]).
Suppose ${\mathbf{A}}$ is a finite algebra, $\rho\in\operatorname{Con}({\mathbf{A}})$ is meet-irreducible, and $\rho^{+}$ is its unique upper cover.
  1. (1)
    If $\operatorname{typ}(\rho,\rho^{+})\in\{\mathbf{2},\mathbf{3}\}$, then $\overline{\rho}\leq{\mathbf{A}}^{2}$ and $\operatorname{Cov}(\rho)=\{\overline{\rho}\}$.
  2. (2)
    If $\operatorname{typ}(\rho,\rho^{+})\in\{\mathbf{4},\mathbf{5}\}$, then $|\operatorname{Cov}(\rho)|=2$, say $\operatorname{Cov}(\rho)=\{\tau_{0},\tau_{1}\}$. Moreover, $\tau_{1}=\tau_{0}^{-1}$, $\tau_{0}\cap\tau_{1}=\rho$, and $\tau_{0}\cup\tau_{1}=\overline{\rho}$.
Now we can characterize irreducible congruences in finite Taylor algebras.
Lemma 5.5.
Suppose ${\mathbf{A}}$ is a finite Taylor algebra, $\rho\in\operatorname{Con}({\mathbf{A}})$ is meet-irreducible, and $\rho^{+}$ is its unique upper cover.
  1. (1)
    The following are equivalent:
    1. (a)
      $\rho$ is irreducible.
    2. (b)
      $\operatorname{typ}(\rho,\rho^{+})\in\{\mathbf{2},\mathbf{3}\}$, and for all $(a,b)\in A^{2}\setminus\rho^{+}$ there exists a unary polynomial $f(x)\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $(f(a),f(b))\in\rho^{+}\setminus\rho$.
  2. (2)
    If $\rho$ is irreducible, then $\rho^{\ast}=\overline{\rho}$.
Proof.
Assume $\rho$ is irreducible. We have $\operatorname{typ}(\rho,\rho^{+})\neq\mathbf{1}$ by Theorem 2.9. Irreducibility of $\rho$ forces $|\operatorname{Cov}(\rho)|=1$, so $\operatorname{typ}(\rho,\rho^{+})\in\{\mathbf{2},\mathbf{3}\}$ by Proposition 5.4(2). Assume next that there exists $(a,b)\in A^{2}\setminus\rho^{+}$ such that $(f(a),f(b))\not\in\rho^{+}\setminus\rho$ for all $f\in\operatorname{Pol}_{1}({\mathbf{B}})$. Let $\sigma=\operatorname{Sg}^{{\mathbf{A}}^{2}}(\{(a,b)\}\cup 0_{A})$. The assumption implies $\sigma\cap\rho^{+}\subseteq\rho$, so $(\rho\circ\sigma\circ\rho)\cap\rho^{+}=\rho$, which would contradict irreducibility of $\rho$. Thus if $\rho$ is irreducible, then the conditions in item (1b) hold.
Conversely, assume that the conditions in item (1b) hold. Since $\operatorname{typ}(\rho,\rho^{+})\in\{\mathbf{2},\mathbf{3}\}$, we have $\operatorname{Cov}(\rho)=\{\overline{\rho}\}$ by Proposition 5.4(1). We will show that $\rho$ is irreducible with $\rho^{\ast}=\overline{\rho}$, which will also establish item (2). Let $R$ be a $\rho$-saturated subuniverse of ${\mathbf{A}}^{2}$ which properly contains $\rho$; we must show $\overline{\rho}\subseteq R$. It will suffice to prove $R\cap\rho^{+}\neq\rho$, as then $R\cap\rho^{+}$ will be a $\rho$-saturated subuniverse of ${\mathbf{A}}^{2}$ satisfying $\rho\subset R\cap\rho^{+}\subseteq\rho^{+}$, so $\overline{\rho}\subseteq R\cap\rho^{+}$ as $\operatorname{Cov}(\rho)=\{\overline{\rho}\}$ by Proposition 5.4(1). To prove $R\cap\rho^{+}\neq\rho$, pick $(a,b)\in R\setminus\rho$. If $(a,b)\in\rho^{+}$ then we are done, so assume $(a,b)\not\in\rho^{+}$. By condition (1b), there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $(a^{\prime},b^{\prime}):=(f(a),f(b))\in\rho^{+}\setminus\rho$. Then $(a^{\prime},b^{\prime})\in\operatorname{Sg}^{{\mathbf{A}}^{2}}(\{(a,b)\}\cup 0_{A})\subseteq R$, so $(a^{\prime},b^{\prime})$ witnesses $R\cap\rho^{+}\neq\rho$. ∎
The following easy lemma will help us extend the notions of $\operatorname{Opt}(\rho)$ and adjacency from irreducible congruences to meet-irreducible congruences.
Lemma 5.6.
Suppose ${\mathbf{A}}$ is a finite Taylor algebra, $\rho\in\operatorname{Con}({\mathbf{A}})$ is meet-irreducible, and $\tau,\tau^{\prime}\in\operatorname{Cov}(\rho)$. Then there exists a reflexive bridge from $({\mathbf{A}},\rho,\tau)$ to $({\mathbf{A}},\rho,\tau^{\prime})$. Hence $\operatorname{Opt}(\rho,\tau)=\operatorname{Opt}(\rho,\tau^{\prime})$.
Proof.
If $\tau^{\prime}=\tau$, then we can use the identity bridge $I_{({\mathbf{A}},\rho,\tau)}$. Otherwise, by Proposition 5.4 we must have $\operatorname{typ}(\rho,\rho^{+})\in\{\mathbf{4},\mathbf{5}\}$ and $\tau^{\prime}=\tau^{-1}$. Then the set $T=\{(a_{1},a_{2},b_{1},b_{2})\,:\,\text{$(a_{1},a_{2},b_{2},b_{1})\in I_{({\mathbf{A}},\rho,\tau)}$}\}$ is the required reflexive bridge. The last claim follows by Lemma 4.16. ∎
Proposition 5.4(1) and Lemmas 5.5(2) and 5.6 justify the the following extension of the notation $\operatorname{Opt}(\rho)$ from irreducible congruences to meet-irreducible congruences.
Definition 5.7.
Suppose ${\mathbf{A}}$ is a finite Taylor algebra and $\rho\in\operatorname{Con}({\mathbf{A}})$ is meet-irreducible.
  1. (1)
    $\operatorname{Opt}(\rho)$ denotes the (unique) congruence $\operatorname{Opt}(\rho,\tau)$ where $\tau\in\operatorname{Cov}(\rho)$.
  2. (2)
    A bridge $T$ from $({\mathbf{A}},\rho)$ to $({\mathbf{A}},\rho)$ is optimal if it is rooted and $\operatorname{tr}(T)=\operatorname{Opt}(\rho)$.
We also extend the adjacency relation to meet-irreducible congruences, as follows.
Definition 5.8.
Suppose ${\mathbf{A}}$ is a finite Taylor algebra and $\rho,\sigma\in\operatorname{Con}({\mathbf{A}})$ are meet-irreducible. Say that $\rho$ and $\sigma$ are adjacent if there exists a reflexive rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{A}},\sigma)$.
Remark 5.9.
Our definition of adjacency in the above context differs from Zhuk’s definition in [18]; Zhuk did not explicitly require the bridge to be rooted. (Our definitions do agree when restricted to irreducible congruences.)
Corollary 4.18 extends to meet-irreducible congruences. The proof is a simple application of Lemmas 4.16 and 5.6.
Corollary 5.10.
Suppose ${\mathbf{A}}$ is a finite Taylor algebra and $\rho,\sigma\in\operatorname{Con}({\mathbf{A}})$ are meet-irreducible. If $\rho$ and $\sigma$ are adjacent, then $\operatorname{Opt}(\rho)=\operatorname{Opt}(\sigma)$.
6. Connecting bridges to centrality and similarity
In this final section we give our main results. We prove that, in the context developed in the previous section, the $\operatorname{Opt}(\rho)$ construction is simply the centralizer $(\rho:\rho^{+})$ (Lemma 6.2); we apply this and results from [16] to extend two important results from [18] (Theorem 6.4 and Lemma 6.7); and we prove that the “there exists a rooted bridge” relation between pairs $({\mathbf{A}},\rho)$ and $({\mathbf{B}},\sigma)$ is exactly the similarity relation between the respective quotient algebras ${\mathbf{A}}/\rho$ and ${\mathbf{B}}/\sigma$ (Corollary 6.11).
Definition 6.1.
Suppose ${\mathbf{A}}$ is a finite Taylor algebra and $\rho\in\operatorname{Con}({\mathbf{A}})$ is meet-irreducible with unique upper cover $\rho^{+}$. If $\rho^{+}/\rho$ is abelian, let $\alpha=(\rho:\rho^{+})$, define
$\Delta^{\flat}_{\rho^{+},\alpha}=\{(a_{1},a_{2},b_{1},b_{2})\in A^{4}\,:\,\text{$((a_{1},a_{2}),(b_{1},b_{2}))\in\Delta_{\rho^{+},\alpha}$}\},$
and set $T^{\,\mathsf{opt}}_{({\mathbf{A}},\rho)}=I_{({\mathbf{A}},\rho,\rho^{+})}\circ\Delta_{\rho^{+},\alpha}^{\flat}\circ I_{({\mathbf{A}},\rho,\rho^{+})}$.
Lemma 6.2.
Suppose ${\mathbf{A}}$ is a finite Taylor algebra and $\rho\in\operatorname{Con}({\mathbf{A}})$ is meet-irreducible with unique upper cover $\rho^{+}$.
  1. (1)
    $\operatorname{Opt}(\rho)=(\rho:\rho^{+})$.
  2. (2)
    Hence $\rho^{+}/\rho$ is abelian if and only if $\operatorname{Opt}(\rho)>\rho$.
  3. (3)
    If $\rho^{+}/\rho$ is abelian, then $\overline{\rho}=\rho^{+}$, so $\operatorname{Cov}(\rho)=\{\rho^{+}\}$, and $T^{\,\mathsf{opt}}_{({\mathbf{A}},\rho)}$ is an optimal bridge from $({\mathbf{A}},\rho)$ to itself.
Proof.
We first show if $\rho^{+}/\rho$ is abelian, i.e., $\operatorname{typ}(\rho,\rho^{+})=\mathbf{2}$, then $\overline{\rho}=\rho^{+}$. Note that $\overline{\rho}/\rho$ is a reflexive subuniverse of $({\mathbf{A}}/\rho)^{2}$ contained in the abelian minimal congruence $\rho^{+}/\rho$ and properly containing $0_{A/\rho}$. ${\mathbf{A}}/\rho$ has a weak difference term by Theorem 2.9, which is a Maltsev operation when restricted to each block of $\rho^{+}/\rho$. Hence $\overline{\rho}/\rho=\rho^{+}/\rho$ by Lemma 2.6. Since $\overline{\rho}$ is $\rho$-saturated, it follows that $\overline{\rho}=\rho^{+}$. Hence $\operatorname{Cov}(\rho)=\{\rho^{+}\}$ by Proposition 5.4(1).
Next we show that if $\rho^{+}/\rho$ is abelian, then $T^{\,\mathsf{opt}}_{({\mathbf{A}},\rho)}$ is a bridge from $({\mathbf{A}},\rho,\rho^{+})$ to itself with trace $\alpha:=(\rho:\rho^{+})$. Note that $\Delta^{\flat}_{\rho^{+},\alpha}$ is not necessarily a bridge from $({\mathbf{A}},\rho,\rho^{+})$ to itself, because it may fail to satisfy (B0). However it does satisfy (B1) and (B2), the latter by Lemma 3.2 using $C(\alpha,\rho^{+};\rho)$. Pre- and post-composing $\Delta^{\flat}_{\rho^{+},\alpha}$ with $I_{({\mathbf{A}},\rho,\rho^{+})}$ preserves (B1) and (B2) and also guarantees (B0). Thus $T^{\,\mathsf{opt}}_{({\mathbf{A}},\rho)}$ is indeed a bridge from $({\mathbf{A}},\rho,\rho^{+})$ to itself. Finally,
$\operatorname{tr}(T^{\,\mathsf{opt}}_{({\mathbf{A}},\rho)})=\operatorname{tr}(I_{({\mathbf{A}},\rho,\rho^{+})})\circ\operatorname{tr}(\Delta^{\flat}_{\rho^{+},\alpha})\circ\operatorname{tr}(I_{({\mathbf{A}},\rho,\rho^{+})})=\rho\circ\alpha\circ\rho=\alpha$
as required. Clearly $T^{\,\mathsf{opt}}_{({\mathbf{A}},\rho)}$ is rooted.
Next, we will prove (1). Assume first that $\rho^{+}/\rho$ is nonabelian. Because $\rho$ is meet-irreducible and $\rho^{+}$ is its unique upper cover, we get $(\rho:\rho^{+})=\rho$. As $\operatorname{Opt}(\rho)\geq\rho$ by Lemma 4.13, it will be enough in this case to show that $\operatorname{Opt}(\rho)\ngeq\rho^{+}$. Assume instead that $\operatorname{Opt}(\rho)\geq\rho^{+}$. Choose $\tau\in\operatorname{Cov}(\rho)$ and let $T_{0}$ be an optimal bridge from $({\mathbf{A}},\rho,\tau)$ to itself. Let $T=T_{0}\circ T_{0}$; then $T$ is also an optimal bridge from $({\mathbf{A}},\rho,\tau)$ to itself. Hence $\operatorname{tr}(T)=\operatorname{Opt}(\rho)$, and moreover $(a,b,a,b)\in T$ for all $(a,b)\in\tau$.
Pick a $(\rho,\rho^{+})$-trace $N$. By tame congruence theory, i.e., Proposition 2.8, there exists $(0,1)\in N^{2}\setminus\rho$ and a binary polynomial $p(x,y)\in\operatorname{Pol}_{2}({\mathbf{A}})$ such that $p(0,0)=p(0,1)=p(1,0)=0$ and $p(1,1)=1$. Then $(0,1)\in\overline{\rho}$ (see Definition 5.3), and since $\overline{\rho}=\tau\cup\tau^{-1}$ by Proposition 5.4, we have either $(0,1)\in\tau$ or $(1,0)\in\tau$. Assume with no loss of generality that $(0,1)\in\tau$.
Choose a $(2+n)$-ary term and a tuple $\mathbf{c}\in A^{n}$ so that $p(x,y)=t(x,y,\mathbf{c})$. Then we have the following tuples in $T$:
[…]
Applying $t$ coordinate-wise to these tuples gives $(0,0,0,1)\in T$, contradicting (B2) since $(0,1)\not\in\rho$. This contradiction proves $\operatorname{Opt}(\rho)\ngeq\rho^{+}$ and hence $\operatorname{Opt}(\rho)=\rho$ when $\rho^{+}/\rho$ is nonabelian.
Assume next that $\rho^{+}/\rho$ is abelian. We will first show $\operatorname{Opt}(\rho)\leq(\rho:\rho^{+})$. As shown earlier, $\operatorname{Cov}(\rho)=\{\rho^{+}\}$. Fix an optimal bridge $T$ from $({\mathbf{A}},\rho,\rho^{+})$ to itself. By replacing $T$ with $T\circ T$, we may assume that $(a,b,a,b)\in T$ for all $(a,b)\in\rho^{+}$. To prove $\operatorname{Opt}(\rho)\leq(\rho:\rho^{+})$, we will simply show that $C(\operatorname{tr}(T),\rho^{+};\rho)$ holds by verifying the condition in Definition 2.2(3).
Let $t(x,\mathbf{y})$ be a $(1+n)$-ary term, let $(a,b)\in\operatorname{tr}(T)$, and let $(c_{i},d_{i})\in\rho^{+}$ for $i=1,\ldots,n$. Then we have the following tuples in $T$:
$(~{}a,~{}a,~{}b,~{}b~{})\in T$
$(c_{1},d_{1},c_{1},d_{1})\in T$
$\vdots$
$(c_{n},d_{n},c_{n},d_{n})\in T$
Applying $t$ coordinatewise gives
$(t(a,\mathbf{c}),t(a,\mathbf{d}),t(b,\mathbf{c}),t(b,\mathbf{d}))\in T.$
Then by (B2),
$t(a,\mathbf{c})\stackrel{{\scriptstyle\rho}}{{\equiv}}t(a,\mathbf{d})\iff t(b,\mathbf{c})\stackrel{{\scriptstyle\rho}}{{\equiv}}t(b,\mathbf{d}).$
This proves that $C(\operatorname{tr}(T),\rho^{+};\rho)$ holds and hence $\operatorname{Opt}(\rho)=\operatorname{tr}(T)\leq(\rho:\rho^{+})$.
On the other hand, we have already shown that $T^{\,\mathsf{opt}}_{({\mathbf{A}},\rho)}$ is a rooted bridge from $({\mathbf{A}},\rho)$ to itself with trace $(\rho:\rho^{+})$. Hence $(\rho:\rho^{+})\leq\operatorname{Opt}(\rho)$, which proves $\operatorname{Opt}(\rho)=(\rho:\rho^{+})$ in the abelian case and completes the proof of (1) and (3). (2) follows from (1). ∎
Using our results about similarity, we can now easily obtain (and generalize) one of the key results in Zhuk [18].
Definition 6.3.
An algebra is affine if it is abelian and has a Maltsev term.
Theorem 6.4 (Cf. Zhuk [18, Corollary 8.17.1]).
Suppose ${\mathbf{A}}$ is a finite Taylor algebra, $\rho\in\operatorname{Con}({\mathbf{A}})$ is irreducible, and $\operatorname{Opt}(\rho)=1_{A}$. Then there exists a simple affine algebra ${\mathbf{Z}}\in\mathsf{HS}({\mathbf{A}}^{2})$ having a 1-element subuniverse $\{0\}\leq{\mathbf{Z}}$, and there exists a subdirect subuniverse $\zeta\leq_{sd}{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{Z}}$ with $\operatorname{pr}_{1,2}(\zeta)=\rho^{\ast}$, such that for all $(a,a^{\prime},b)\in\zeta$,
$(a,a^{\prime})\in\rho\iff b=0.$
Proof.
Let $\rho^{+}$ be the unique upper cover of $\rho$. By Lemma 6.2, we have $(\rho:\rho^{+})=1_{A}$ and $\rho^{+}/\rho$ is abelian. Thus $\rho^{\ast}=\rho^{+}$ by Lemmas 5.5(2) and 6.2(3). Let $\overline{{\mathbf{A}}}={\mathbf{A}}/\rho$ and $\mu=\rho^{+}/\rho$; thus $\overline{{\mathbf{A}}}$ is subdirectly irreducible with abelian monolith $\mu$ satisfying $(0:\mu)=1$. Let ${\mathbf{Z}}=D(\overline{{\mathbf{A}}})$. By Theorem 3.4, ${\mathbf{Z}}$ is simple and abelian, so is affine, and has a 1-element subuniverse $D^{o}=\{0\}$. Also by Theorem 3.4, there exists a surjective homomorphism $h:\overline{{\mathbf{A}}}(\mu)\rightarrow{\mathbf{Z}}$ such that $h^{-1}(0)=0_{\overline{A}}$. Let
$\zeta=\{(a,a^{\prime},b)\in A\times A\times Z\,:\,\text{$(a,a^{\prime})\in\rho^{+}$ and $h((a/\rho,a^{\prime}/\rho))=b$}\}.$
$\zeta$ has the required properties. ∎
Remark 6.5.
Zhuk [18] proved Theorem 6.4 in the special case where the signature of ${\mathbf{A}}$ consists of just one operation, $w(x_{1},\ldots,x_{m})$, which is an $m$-ary special WNU, that is, an (idempotent) weak near-unanimity operation whose derived binary operation $x\circ y:=w(x,\ldots,x,y)$ satisfies $x\circ(x\circ y)=x\circ y$. In this context, if ${\mathbf{Z}}\in\mathsf{HS}({\mathbf{A}}^{2})$ then $w$ is also an $m$-ary special WNU in ${\mathbf{Z}}$; if in addition ${\mathbf{Z}}$ is simple and affine, then it is not hard to show (cf. [19, Lemma 6.4]) that ${\mathbf{Z}}\cong(\mathbb{Z}_{p},x_{1}+\cdots+x_{m}\pmod{p})$ for some prime $p$ which is a divisor of $m-1$. Zhuk stated his [18, Corollary 8.17.1] with this stronger conclusion.
Next we establish a simple invariant of rooted bridges.
Theorem 6.6.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are finite algebras in a locally finite Taylor variety, and $\rho\in\operatorname{Con}({\mathbf{A}})$, $\sigma\in\operatorname{Con}({\mathbf{B}})$, where $\rho$ and $\sigma$ are meet-irreducible. Let $\rho^{+},\sigma^{+}$ be the respective unique upper covers of $\rho,\sigma$. Assume that there exists a rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$. Then $\rho^{+}/\rho$ is abelian if and only if $\sigma^{+}/\sigma$ is abelian.
Proof.
Let $T$ be a rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$. We may assume that $T$ is from $({\mathbf{A}},\rho,\tau)$ to $({\mathbf{B}},\sigma,\tau^{\prime})$ where $\tau\in\operatorname{Cov}(\rho)$ and $\tau^{\prime}\in\operatorname{Cov}(\sigma)$. Assume for the sake of contradiction that $\rho^{+}/\rho$ is abelian while $\sigma^{+}/\sigma$ is not. Then by tame congruence theory, $\operatorname{typ}(\rho,\rho^{+})=\mathbf{2}$ while $\operatorname{typ}(\sigma,\sigma^{+})\in\{\mathbf{3},\mathbf{4},\mathbf{5}\}$. Hence $\tau=\rho^{+}$ by Lemma 6.2 and $\tau^{\prime}\cup(\tau^{\prime})^{-1}=\overline{\sigma}$ by Proposition 5.4.
By passing to $\overline{{\mathbf{A}}}:={\mathbf{A}}/\rho$ and $\overline{{\mathbf{B}}}:={\mathbf{B}}/\sigma$, we may assume that $\rho=0_{A}$ and $\sigma=0_{B}$. Rename $\rho^{+}$ and $\sigma^{+}$ as $\mu_{{\mathbf{A}}}$ and $\mu_{{\mathbf{B}}}$ respectively. Let $\alpha=(0_{A}:\mu_{{\mathbf{A}}})$ and observe that $(0_{B}:\mu_{{\mathbf{B}}})=0_{B}$. Replace $T$ with $T^{\,\mathsf{opt}}_{({\mathbf{A}},0)}\circ T$; then by Lemma 4.15, the rule
$h(a)=b\iff(a,b)\in\operatorname{tr}(T)$
defines a surjective homomorphism $h:{\mathbf{A}}\rightarrow{\mathbf{B}}$ with kernel $\alpha$.
Claim 6.6.1.
For all $f\in\operatorname{Pol}_{k}({\mathbf{B}})$ there exists $f_{{\mathbf{A}}}\in\operatorname{Pol}_{k}({\mathbf{A}})$ such that
  1. (1)
    $h(f_{{\mathbf{A}}}(x_{1},\ldots,x_{k}))=f(h(x_{1}),\ldots,h(x_{k}))$ for all $x_{1},\ldots,x_{k}\in A$.
  2. (2)
    For all $(x_{1},y_{1},u_{1},v_{1}),\ldots,(x_{k},y_{k},u_{k},v_{k})\in T$ we have
    $(f_{{\mathbf{A}}}(\mathbf{x}),f_{{\mathbf{A}}}(\mathbf{y}),f(\mathbf{u}),f(\mathbf{v}))\in T.$
Proof of Claim 6.6.1.
Indeed, if we select a term $t(x_{1},\ldots,x_{k},y_{1},\ldots,y_{n})$ and $\mathbf{b}\in B^{n}$ so that $f(\mathbf{x})=t^{{\mathbf{B}}}(\mathbf{x},\mathbf{b})$, then we simply need to select $\mathbf{a}\in A^{n}$ with $h(a_{i})=b_{i}$ for each $i\in[n]$ and then define $f_{{\mathbf{A}}}(\mathbf{x})=t^{{\mathbf{A}}}(\mathbf{x},\mathbf{a})$. Item (1) then follows immediately, and item (2) follows from the fact that $(a_{i},a_{i},b_{i},b_{i})\in T$ for all $i\in[n]$. ∎
By tame congruence theory, i.e., Proposition 2.8, there exists a $(0_{B},\mu_{{\mathbf{B}}})$-minimal set $U$ with unique $(0_{B},\mu_{{\mathbf{B}}})$-trace $N=\{0,1\}=U\cap C$, a unary polynomial $e(x)\in\operatorname{Pol}_{1}({\mathbf{B}})$, and a binary polynomial $p(x,y)\in\operatorname{Pol}_{2}({\mathbf{B}})$, satisfying:
  1. (1)
    $e(A)=U$ and $e(e(x))=e(x)$ for all $x\in A$.
  2. (2)
    $p(x,x)=p(x,1)=p(1,x)=x$ for all $x\in U$.
  3. (3)
    $p(x,0)=p(0,x)=x$ for all $x\in U\setminus\{1\}$.
Because $T$ is rooted, we have $(0,1)\in\operatorname{pr}_{3,4}(T)$ or $(1,0)\in\operatorname{pr}_{3,4}(T)$. Assume with no loss of generality that $(0,1)\in\operatorname{pr}_{3,4}(T)$. Pick $(a^{\prime},b^{\prime})\in\mu_{{\mathbf{A}}}$ with $(a^{\prime},b^{\prime},0,1)\in T$. Let $e_{{\mathbf{A}}}\in\operatorname{Pol}_{1}({\mathbf{A}})$ be a polynomial given by Claim 6.6.1 for $e(x)$. Let $a=e_{{\mathbf{A}}}(a^{\prime})$ and $b=e_{{\mathbf{A}}}(b^{\prime})$. Then $(a,b,0,1)\in T$ and $h(a)=e(h(a^{\prime}))\in U$ by Claim 6.6.1. Since $(a,b)\in\mu_{{\mathbf{A}}}\subseteq\ker(h)$ we get $h(b)=h(a)$. Let $u=h(a)\in U$.
Now let $p_{{\mathbf{A}}}\in\operatorname{Pol}_{2}({\mathbf{A}})$ be a polynomial given by Claim 6.6.1 for $p(x,y)$. In calculations that follow, we will denote both $p(x,y)$ and $p_{{\mathbf{A}}}(x,y)$ by $xy$.
We have the following elements of $T$:
$\tau_{1}=(x_{1},y_{1},u_{1},v_{1}):={}$ $(a,b,0,1)$
$\tau_{2}=(x_{2},y_{2},u_{2},v_{2}):={}$ $(a,a,u,u)$
$\tau_{3}=(x_{3},y_{3},u_{3},v_{3}):={}$ $(b,b,u,u).$
Case 1: $u\neq 1$.
Applying Claim 6.6.1(2) to $p(x,y)$ and the pairs $(\tau_{1},\tau_{1})$, $(\tau_{2},\tau_{1})$ and $(\tau_{1},\tau_{3})$ respectively gives
(6.1) $(aa,bb,0,1)$ $\in T$
(6.2) $(aa,ab,u,u)$ $\in T$
(6.3) $(ab,bb,u,u)$ $\in T.$
(6.1) with (B2) gives $aa\neq bb$, but (6.2) and (6.3) with (B2) give $aa=ab=bb$, contradiction.
Case 2: $u=1$.
Applying Claim 6.6.1(2) to $p(x,y)$ and the pairs $(\tau_{1},\tau_{1})$, $(\tau_{2},\tau_{1})$ and $(\tau_{1},\tau_{3})$ respectively gives
(6.4) $\sigma_{1}:=(aa,bb,0,1)$ $\in T$
(6.5) $\sigma_{2}:=(aa,ab,0,1)$ $\in T$
(6.6) $\sigma_{3}:=(ab,bb,0,1)$ $\in T.$
Let $d(x,y,z)$ be a weak difference term for the locally finite Taylor variety containing ${\mathbf{A}}$ and ${\mathbf{B}}$. Recall that $\mu_{{\mathbf{A}}}$ is abelian and observe that $aa,ab,bb$ all belong to a common $\mu_{{\mathbf{A}}}$-class. Applying $d$ coordinate-wise to the tuples $\sigma_{2},\sigma_{1},\sigma_{3}$ (in that order) and using the defining property of weak difference terms, we get $(ab,ab,0,1)\in T$, which again contradicts (B2).
As we have found a contradiction in both cases, the theorem is proved. ∎
Now we can extend and give a relatively short proof of an important result about adjacent congruences in [18].
Lemma 6.7 (cf. Zhuk [18, Lemma 8.18]).
Suppose ${\mathbf{A}}$ is a finite Taylor algebra and $\rho,\sigma\in\operatorname{Con}({\mathbf{A}})$ are meet-irreducible and adjacent. If $\rho\neq\sigma$, then $\operatorname{Opt}(\sigma)>\sigma$.
Proof.
Let $\rho^{+},\sigma^{+}$ be the unique upper covers of $\rho,\sigma$ respectively. Assume $\operatorname{Opt}(\sigma)=\sigma$. Then $\sigma^{+}/\sigma$ is nonabelian by Lemma 6.2(2). Hence $\rho^{+}/\rho$ is nonabelian by Theorem 6.6, so $\rho=\operatorname{Opt}(\rho)=\operatorname{Opt}(\sigma)=\sigma$ by Lemma 6.2(2) and Corollary 4.18, contradicting $\rho\neq\sigma$. ∎
For the remainder of this section, we work to characterize the “there exists a rooted bridge” relation between meet-irreducible congruences. The next lemma handles the nonabelian case.
Lemma 6.8.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are finite algebras in a locally finite Taylor variety and $\rho\in\operatorname{Con}({\mathbf{A}})$, $\sigma\in\operatorname{Con}({\mathbf{B}})$, where $\rho$ and $\sigma$ are meet-irreducible. Let $\rho^{+},\sigma^{+}$ be the respective unique upper covers of $\rho,\sigma$.
  1. (1)
    If ${\mathbf{A}}/\rho\cong{\mathbf{B}}/\sigma$, then there exists a rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$.
  2. (2)
    Conversely, if $\rho^{+}/\rho$ and $\sigma^{+}/\sigma$ are both nonabelian and there exists a rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$, then ${\mathbf{A}}/\rho\cong{\mathbf{B}}/\sigma$.
Proof.
(1) Suppose $\gamma:{\mathbf{A}}/\rho\cong{\mathbf{B}}/\sigma$ is an isomorphism. Then
$T=\{(a,a^{\prime},b,b^{\prime})\in A\times A\times B\times B\,:\,\text{$\gamma(a/\rho)=b/\sigma$ and $\gamma(a^{\prime}/\rho)=b^{\prime}/\sigma$}\}$
is a rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$.
(2) We have $\operatorname{Opt}(\rho)=\rho$ and $\operatorname{Opt}(\sigma)=\sigma$ by Lemma 6.2(2). Let $T$ be a rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$. We can assume that $T$ is a bridge from $({\mathbf{A}},\rho,\tau)$ to $({\mathbf{B}},\sigma,\tau^{\prime})$ for some $\tau\in\operatorname{Cov}(\rho)$ and $\tau^{\prime}\in\operatorname{Cov}(\sigma)$. Thus $\rho=\operatorname{Opt}(\rho,\tau)$ and $\sigma=\operatorname{Opt}(\sigma,\tau^{\prime})$. Now the claim follows from Lemma 4.15. ∎
It remains to characterize the “there exists a rooted bridge” relation between meet-irreducible congruences $\rho$ and $\sigma$ when $\rho^{+}/\rho$ and $\sigma^{+}/\sigma$ are both abelian. We will see that there is a tight relationship to similarity and proper bridges as defined in [16]. The main difficulty is that, although Zhuk’s definition of bridges is similar to the definition of proper bridges in [16], the definitions differ in one essential way: we required proper bridges to satisfy
  1. (B3)
    For all $(a_{1},a_{2},b_{1},b_{2})\in T$ we have $(a_{i},b_{i})\in\operatorname{tr}(T)$ for $i=1,2$,
while Zhuk’s bridges are not required to satisfy this condition. Happily, rooted bridges between meet-irreducible congruences with abelian upper covers can be assumed without loss of generality to satisfy (B3), as we will prove in Theorem 6.10. First, we need the following result about the optimal bridges from Lemma 6.2(3).
Lemma 6.9.
Suppose ${\mathbf{A}}$ is a finite subdirectly Taylor algebra with abelian monolith $\mu$. Let $\alpha=(0:\mu)$ and define $\Delta^{\flat}_{\mu,\alpha}$ as in Definition 6.1 (setting $\rho:=0$, so $\rho^{+}=\mu$). Recall the proper bridge $T_{{\mathbf{A}}}^{D}$ from ${\mathbf{A}}$ to $D({\mathbf{A}})$ defined in Corollary 3.9. Then $T^{\,\mathsf{opt}}_{({\mathbf{A}},0)}=\Delta^{\flat}_{\mu,\alpha}=T_{{\mathbf{A}}}^{D}\circ(T_{{\mathbf{A}}}^{D})^{\cup}$.
Proof.
The first equality follows from the fact that $I_{({\mathbf{A}},0,\mu)}=\{(a,b,a,b)\,:\,\text{$(a,b)\in\mu$}\}$. Let $\Delta=\Delta_{\mu,\alpha}$. The second equality will follow if we can show that for all $(a_{1},a_{2})$, $(b_{1},b_{2})\in\mu$ with $a_{1},a_{2},b_{1},b_{2}$ belonging to a common $\alpha$-class,
$(a_{1},a_{2})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b_{1},b_{2})\iff\exists e\stackrel{{\scriptstyle\mu}}{{\equiv}}a_{1},\exists e^{\prime}\stackrel{{\scriptstyle\mu}}{{\equiv}}b_{1}\left((a_{1},e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b_{1},e^{\prime})~{}\&~{}(a_{2},e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b_{2},e^{\prime})\right).$
The forward implication is easy: choose $e=a_{2}$ and $e^{\prime}=b_{2}$. For the reverse implication, apply a weak difference term component-wise to
$(a_{1},e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b_{1},e^{\prime}),\quad(a_{2},e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b_{2},e^{\prime}),\quad(a_{2},a_{2})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b_{2},b_{2})$
to get $(a_{1},a_{2})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(b_{1},b_{2})$. ∎
Theorem 6.10.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are finite algebras in a locally finite Taylor variety, $\rho\in\operatorname{Con}({\mathbf{A}})$ and $\sigma\in\operatorname{Con}({\mathbf{B}})$ where both $\rho$ and $\sigma$ are meet-irreducible, and $\rho^{+},\sigma^{+}$ are their respective unique upper covers. Assume that $\rho^{+}/\rho$ and $\sigma^{+}/\sigma$ are abelian. Then for every bridge $T$ from $({\mathbf{A}},\rho,\rho^{+})$ to $({\mathbf{B}},\sigma,\sigma^{+})$, the bridge $T^{\prime}:=T^{\,\mathsf{opt}}_{({\mathbf{A}},\rho)}\circ T\circ T^{\,\mathsf{opt}}_{({\mathbf{B}},\sigma)}$ contains a bridge from $({\mathbf{A}},\rho,\rho^{+})$ to $({\mathbf{B}},\sigma,\sigma^{+})$ with the same trace as $T^{\prime}$ and satisfying (B3).
Proof.
First, we can assume with no loss of generality that $T^{\prime}=T$. Second, we can assume that $\rho=0_{A}$ and $\sigma=0_{B}$. For we can let $\overline{{\mathbf{A}}}={\mathbf{A}}/\rho$, $\overline{{\mathbf{B}}}={\mathbf{B}}/\sigma$, $\mu=\rho^{+}/\rho$, $\kappa=\sigma^{+}/\sigma$, and
$\overline{T}=\{(a_{1}/\rho,a_{2}/\rho,b_{1}/\sigma,b_{2}/\sigma)\,:\,\text{$(a_{1},a_{2},b_{1},b_{2})\in T$}\}$
and $\overline{T}$ will be a bridge from $(\overline{{\mathbf{A}}},0,\mu)$ to $(\overline{{\mathbf{B}}},0,\kappa)$ satisfying $\overline{T}=T_{(\overline{{\mathbf{A}}},0)}\circ\overline{T}\circ T_{(\overline{{\mathbf{B}}},0)}$. If there exists a bridge $T_{1}$ from $(\overline{{\mathbf{A}}},0,\mu)$ to $(\overline{{\mathbf{B}}},0,\kappa)$ satisfying $T_{1}\subseteq\overline{T}$, $\operatorname{tr}(T_{1})=\operatorname{tr}(\overline{T})$ and (B3), then $T_{0}:=\{(a_{1},a_{2},b_{1},b_{2})\,:\,\text{$(a_{1}/\rho,a_{2}/\rho,b_{1}/\sigma,b_{2}/\sigma)\in T_{1}$}\}$ will be a bridge from $({\mathbf{A}},\rho,\rho^{+})$ to $({\mathbf{B}},\sigma,\sigma^{+})$ satisfying $T_{0}\subseteq T$, $\operatorname{tr}(T_{0})=\operatorname{tr}(T)$ and (B3).
So for the remainder of this proof assume that $T^{\prime}=T$, $\rho=0_{A}$, and $\sigma=0_{B}$. For readability, rename $\rho^{+}$ as $\mu$ and $\sigma^{+}$ as $\kappa$. Let $\alpha=(0:\mu)$, $\Delta_{{\mathbf{A}}}=\Delta_{\mu,\alpha}$, and $\varphi_{{\mathbf{A}}}=\overline{\alpha}/\Delta_{{\mathbf{A}}}$. Recall from Corollary 3.9 that the set
$T_{{\mathbf{A}}}^{D}=\{(a_{1},a_{2},(a_{1},e)/\Delta_{{\mathbf{A}}},(a_{2},e)/\Delta_{{\mathbf{A}}})\in A\times A\times D({\mathbf{A}})\times D({\mathbf{A}})\,:\,\text{$a_{1}\stackrel{{\scriptstyle\mu}}{{\equiv}}a_{2}\stackrel{{\scriptstyle\mu}}{{\equiv}}e$}\}$
is a proper bridge from ${\mathbf{A}}$ to $D({\mathbf{A}})$, and hence is a bridge from $({\mathbf{A}},0,\mu)$ to $(D({\mathbf{A}}),0,\varphi_{{\mathbf{A}}})$ satisfying (B3).
Similarly define $\beta=(0:\kappa)$, $\Delta_{{\mathbf{B}}}=\Delta_{\kappa,\beta}$, and $\varphi_{{\mathbf{B}}}=\overline{\beta}/\Delta_{{\mathbf{B}}}$; then
$T_{{\mathbf{B}}}^{D}=\{(b_{1},b_{2},(b_{1},u)/\Delta_{{\mathbf{B}}},(b_{2},u)/\Delta_{{\mathbf{B}}})\in B\times B\times D({\mathbf{B}})\times D({\mathbf{B}})\,:\,\text{$b_{1}\stackrel{{\scriptstyle\mu}}{{\equiv}}b_{2}\stackrel{{\scriptstyle\mu}}{{\equiv}}u$}\}$
is a bridge from $({\mathbf{B}},0,\kappa)$ to $(D({\mathbf{B}}),0,\varphi_{{\mathbf{B}}})$ satisfying (B3). Thus by composing, we get the bridge $T^{\ast}:=(T_{{\mathbf{A}}}^{D})^{\cup}\circ T\circ T_{{\mathbf{B}}}^{D}$ from $(D({\mathbf{A}}),0,\varphi_{{\mathbf{A}}})$ to $(D({\mathbf{B}}),0,\varphi_{{\mathbf{B}}})$.
Suppose there exists a bridge $T_{1}^{\ast}$ from $(D({\mathbf{A}}),0,\varphi_{{\mathbf{A}}})$ to $(D({\mathbf{B}}),0,\varphi_{{\mathbf{B}}})$ satisfying $T_{1}^{\ast}\subseteq T^{\ast}$, $\operatorname{tr}(T_{1}^{\ast})=\operatorname{tr}(T^{\ast})$ and (B3). In this case we could define $T_{1}=T_{{\mathbf{A}}}^{D}\circ T_{1}^{\ast}\cup(T_{{\mathbf{B}}}^{D})^{\cup}$. Then $T_{1}$ will be a bridge from $({\mathbf{A}},0,\mu)$ to $({\mathbf{B}},0,\kappa)$. We will have
$T_{1}$ $\subseteq T_{{\mathbf{A}}}^{D}\circ T^{\ast}\circ(T_{{\mathbf{B}}}^{D})^{\cup}$
$=T_{{\mathbf{A}}}^{D}\circ((T_{{\mathbf{A}}}^{D})^{\cup}\circ T\circ T_{{\mathbf{B}}}^{D})\circ(T_{{\mathbf{B}}}^{D})^{\cup}$
$=(T_{{\mathbf{A}}}^{D}\circ(T_{{\mathbf{A}}}^{D})^{\cup})\circ T\circ(T_{{\mathbf{B}}}^{D}\circ(T_{{\mathbf{B}}}^{D})^{\cup})$
$=T^{\,\mathsf{opt}}_{({\mathbf{A}},0)}\circ T\circ T^{\,\mathsf{opt}}_{({\mathbf{B}},0)}$ by Lemma 6.9
$=T$ as we’ve assumed $T^{\prime}=T$.
Hence $\operatorname{tr}(T_{1})\subseteq\operatorname{tr}(T)$. Similarly,
$\operatorname{tr}(T_{1})$ $=\operatorname{tr}(T_{{\mathbf{A}}}^{D})\circ\operatorname{tr}(T_{1}^{\ast})\circ\operatorname{tr}((T_{{\mathbf{B}}}^{D})^{\cup})$
$=\operatorname{tr}(T_{{\mathbf{A}}}^{D})\circ\operatorname{tr}(T^{\ast})\circ\operatorname{tr}((T_{{\mathbf{B}}}^{D})^{\cup})$ $\operatorname{tr}(T_{1}^{\ast})=\operatorname{tr}(T^{\ast})$
$=\operatorname{tr}(T^{\,\mathsf{opt}}_{({\mathbf{A}},0)})\circ\operatorname{tr}(T)\circ\operatorname{tr}(T^{\,\mathsf{opt}}_{({\mathbf{B}},0)})$
$=\alpha\circ\operatorname{tr}(T)\circ\beta,$
which proves $\operatorname{tr}(T)\subseteq\operatorname{tr}(T_{1})$. Hence $\operatorname{tr}(T_{1})=\operatorname{tr}(T)$. Finally, it is easy to check that $T_{1}$ satisfies (B3), since each of $T_{{\mathbf{A}}}^{D},T_{1}^{\ast},T_{{\mathbf{B}}}^{D}$ satisfies (B3).
The remarks in the previous paragraph serve to further reduce the proof of Theorem 6.10 to the case where ${\mathbf{A}}$ and ${\mathbf{B}}$ are replaced by $D({\mathbf{A}})$ and $D({\mathbf{B}})$ respectively (and $\rho=0_{D(A)}$ and $\sigma=0_{D(B)}$ and $T^{\prime}=T$). Put differently, in proving Theorem 6.10, we can further assume with no loss of generality that ${\mathbf{A}}\cong D({\mathbf{A}}_{1})$ and ${\mathbf{B}}\cong D({\mathbf{B}}_{1})$ for some subdirectly irreducible algebras ${\mathbf{A}}_{1},{\mathbf{B}}_{1}$ with abelian monoliths. It follows from this assumption and Theorem 3.4 that $\alpha=\mu$, $\beta=\kappa$, and there exist subuniverses $S_{{\mathbf{A}}}\leq{\mathbf{A}}$ and $S_{{\mathbf{B}}}\leq{\mathbf{B}}$ which are transversals for $\mu$ and $\kappa$ respectively.
Recall that we are assuming $T=T^{\,\mathsf{opt}}_{({\mathbf{A}},0)}\circ T\circ T^{\,\mathsf{opt}}_{({\mathbf{B}},0)}$. In this context this means $T=\Delta^{\flat}_{\mu,\mu}\circ T\circ\Delta^{\flat}_{\kappa,\kappa}$. Hence $\operatorname{tr}(T)=\operatorname{tr}(\Delta^{\flat}_{\mu,\mu})\circ\operatorname{tr}(T)\circ\operatorname{tr}(\Delta^{\flat}_{\kappa,\kappa})=\mu\circ\operatorname{tr}(T)\circ\kappa$.
Recall from Lemma 4.15 that $\operatorname{tr}(T)$ induces an isomorphism $\gamma:{\mathbf{A}}/\mu\cong{\mathbf{B}}/\kappa$ defined by
$\gamma(a/\mu)=b/\kappa\iff(a,b)\in\mu\circ\operatorname{tr}(T)\circ\kappa=\operatorname{tr}(T).$
We also have the homomorphism $\pi_{{\mathbf{A}}}:{\mathbf{A}}\rightarrow{\mathbf{S}}_{{\mathbf{A}}}$ which sends each $a\in A$ to the unique element of $S_{{\mathbf{A}}}\cap a/\mu$. Likewise we have $\pi_{{\mathbf{B}}}:{\mathbf{B}}\rightarrow{\mathbf{S}}_{{\mathbf{B}}}$. These retractions naturally induce isomorphisms $\overline{\pi}_{{\mathbf{A}}}:{\mathbf{A}}/\mu\cong{\mathbf{S}}_{{\mathbf{A}}}$ and $\overline{\pi}_{{\mathbf{B}}}:{\mathbf{B}}/\kappa\cong{\mathbf{S}}_{{\mathbf{B}}}$ given by $\overline{\pi}_{{\mathbf{A}}}(a/\mu)=\pi_{{\mathbf{A}}}(a)$ and $\overline{\pi}_{{\mathbf{B}}}(b/\kappa)=\pi_{{\mathbf{B}}}(b)$. Thus we get an isomorphism $\delta:{\mathbf{S}}_{{\mathbf{A}}}\cong{\mathbf{S}}_{{\mathbf{B}}}$ given by $\delta=\overline{\pi}_{{\mathbf{B}}}\circ\gamma\circ(\overline{\pi}_{{\mathbf{A}}})^{-1}$. Equivalently,
(6.7) $\operatorname{graph}(\delta)=\operatorname{tr}(T)\cap(S_{{\mathbf{A}}}\times S_{{\mathbf{B}}}).$
Claim 6.10.1.
If there exists $(a_{1},a_{2},b_{1},b_{2})\in T$ with $a_{1}\neq a_{2}$ and $(a_{1},b_{1})\in\operatorname{tr}(T)$, then $T$ contains a bridge $T_{1}$ from $({\mathbf{A}},0,\mu)$ to $({\mathbf{B}},0,\kappa)$ satisfying $\operatorname{tr}(T_{1})=\operatorname{tr}(T)$ and (B3).
Proof of Claim 6.10.1.
Let $(a_{1},a_{2},b_{1},b_{2})\in T$ satisfy $a_{1}\neq b_{1}$ and $(a_{1},b_{1})\in\operatorname{tr}(T)$. Observe that we also have $(a_{2},b_{2})\in\operatorname{tr}(T)$ as $\operatorname{tr}(T)=\mu\circ\operatorname{tr}(T)\circ\kappa$, and $b_{1}\neq b_{2}$ by (B1). Let $T_{1}$ be the subuniverse of ${\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ generated by
$\{(a,a,b,b)\,:\,\text{$(a,b)\in\operatorname{tr}(T)$}\}\cup\{(a_{1},a_{2},b_{1},b_{2})\}.$
Then $T_{1}\subseteq T$. We will verify that $T_{1}$ is a bridge from $({\mathbf{A}},0,\mu)$ to $({\mathbf{B}},0,\kappa)$. Property (B0) is trivially true and (B2) is inherited from $T$, so what must be shown is (B1): that $\operatorname{pr}_{1,2}(T_{1})=\mu$ and $\operatorname{pr}_{3,4}(T_{1})=\kappa$. By construction, $\operatorname{pr}_{1,2}(T_{1})$ is a reflexive subuniverse of $\mu$ properly containing $0_{A}$; hence $\operatorname{pr}_{1,2}(T_{1})=\mu$ by Lemma 2.6. A similar argument gives $\operatorname{pr}_{3,4}(T_{1})=\kappa$. Thus $T_{1}$ is a bridge from $({\mathbf{A}},0,\mu)$ to $({\mathbf{B}},0,\kappa)$ satisfying $T_{1}\subseteq T$, and clearly $\operatorname{tr}(T_{1})=\operatorname{tr}(T)$ by construction.
Now let
$W=\{(a,a^{\prime},b,b^{\prime})\in A\times A\times B\times B\,:\,\text{ $(a,b),(a^{\prime},b^{\prime})\in\operatorname{tr}(T)$}\}.$
Note that $W$ is a subuniverse of ${\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$. As the generators of $T_{1}$ are contained in $W$, we get $T_{1}\subseteq W$, which proves that $T_{1}$ satisfies (B3). ∎
The remainder of the proof of Theorem 6.10 will consist of the construction of a tuple $(a_{1},a_{2},b_{1},b_{2})\in T$ satisfying the hypotheses of Claim 6.10.1. Define $R\leq{\mathbf{A}}\times{\mathbf{B}}$ by
$R=\operatorname{pr}_{1,3}(T\cap(A\times S_{{\mathbf{A}}}\times B\times S_{{\mathbf{B}}})).$
Claim 6.10.2.
  1. (1)
    $\operatorname{graph}(\delta)\subseteq R$. Thus $S_{{\mathbf{A}}}\subseteq\operatorname{pr}_{1}(R)$ and $S_{{\mathbf{B}}}\subseteq\operatorname{pr}_{2}(R)$.
  2. (2)
    For all $(a,b)\in R$ we have $a\in S_{{\mathbf{A}}}\iff b\in S_{{\mathbf{B}}}\iff(a,b)\in\operatorname{graph}(\delta)$.
Proof.
(1) Assume $(a,b)\in\operatorname{graph}(\delta)$. Then $(a,b)\in\operatorname{tr}(T)\cap(S_{{\mathbf{A}}}\times S_{{\mathbf{B}}})$ by (6.7), so $(a,a,b,b)\in T\cap(S_{{\mathbf{A}}}\times S_{{\mathbf{A}}}\times S_{{\mathbf{B}}}\times S_{{\mathbf{B}}})$, proving $(a,b)\in R$.
(2) Suppose we have $(a,b)\in R$ with $a\in S_{{\mathbf{A}}}$. Choose $x\in S_{{\mathbf{A}}}$ and $y\in S_{{\mathbf{B}}}$ with $(a,x,y,b)\in T$. As $(a,x)\in\mu$ and $a,x\in S_{{\mathbf{A}}}$, we get $a=x$, which forces $b=y$ by bridge property (B2). Hence $(a,b)\in\operatorname{tr}(T)\cap(S_{{\mathbf{A}}}\times S_{{\mathbf{B}}})=\operatorname{graph}(\delta)$ by (6.7). ∎
Claim 6.10.3.
$\operatorname{pr}_{1}(R)\neq S_{{\mathbf{A}}}$ and $\operatorname{pr}_{2}(R)\neq S_{{\mathbf{B}}}$.
Proof of Claim 6.10.3.
Pick a $\mu$-class $C$ with $|C|>1$. Let $u$ be the unique element in $C\cap S_{{\mathbf{A}}}$ and pick $a\in C\setminus\{u\}$. As $T$ is a bridge, there exists $(b,c)\in\kappa$ with $b\neq c$ and $(a,u,b,c)\in T$. Let $y=\pi_{{\mathbf{B}}}(b)\in S_{{\mathbf{B}}}$ and let $x=\delta^{-1}(y)\in S_{{\mathbf{A}}}$. Then $(x,y)\in\operatorname{graph}(\delta)\subseteq\operatorname{tr}(T)$, and hence $(x,c)\in\operatorname{tr}(T)$ as well (using $\operatorname{tr}(T)=\operatorname{tr}(T)\circ\kappa$).
Thus the following are elements of $T$:
$(x,x,y,y),\quad(x,x,c,c),\quad(a,u,b,c).$
Applying the weak difference term $d(x,y,z)$ to these elements of $T$ gives
(6.8) $(a^{\prime},u^{\prime},b^{\prime},y)\in T$
where $a^{\prime}=d(x,x,a)$, $u^{\prime}=d(x,x,u)$ and $b^{\prime}=d(y,c,b)$. Observe that $u^{\prime}\in S_{{\mathbf{A}}}$ since $x,u\in S_{{\mathbf{A}}}$. Thus (6.8) implies $(a^{\prime},b^{\prime})\in R$. Also note that $b,c,y$ belong to a common $\kappa$-class, so $b^{\prime}=y-c+b$ calculated in the abelian group $\operatorname{Grp}_{{\mathbf{B}}}(\kappa,y)$ by Lemma 2.5, so $b^{\prime}\neq y$ (since $b\neq c$). Thus $b^{\prime}\not\in S_{{\mathbf{B}}}$, proving $\operatorname{pr}_{2}(R)\neq S_{{\mathbf{B}}}$. A similar proof gives $\operatorname{pr}_{1}(R)\neq S_{{\mathbf{A}}}$. ∎
Claim 6.10.4.
$\operatorname{pr}_{1}(R)=A$ and $\operatorname{pr}_{2}(R)=B$.
Proof of Claim 6.10.4.
By Claim 6.10.2, $S_{{\mathbf{A}}}\subseteq\operatorname{pr}_{1}(R)$. By Claim 6.10.3, we can choose $a_{0}\in\operatorname{pr}_{1}(R)$ with $a_{0}\not\in S_{{\mathbf{A}}}$. Let $a\in A$ be arbitrary. By Lemma 2.7, there exists a term $t(x,\mathbf{y})$ and a tuple $\mathbf{u}$ of elements from $S_{{\mathbf{A}}}$ such that $t(a_{0},\mathbf{u})=a$. As $\operatorname{pr}_{1}(R)$ is a subuniverse, this proves $a\in\operatorname{pr}_{1}(R)$, and as $a$ was arbitrary, we have shown $\operatorname{pr}_{1}(R)=A$. A similar proof gives $\operatorname{pr}_{2}(R)=B$. ∎
Now define
$\theta:=\{(u,u^{\prime})\in(S_{{\mathbf{A}}})^{2}\,:\,\text{$\exists(a,b)\in R$ with $\pi_{{\mathbf{A}}}(a)=u$ and $\pi_{{\mathbf{B}}}(b)=\delta(u^{\prime})$}\}.$
If $b,c$ are two elements from $A$ or two elements from $B$, we will say that $(b,c)$ is a Maltsev pair if $d(b,b,c)=c=d(c,b,b)$.
Claim 6.10.5.
  1. (1)
    For all $(a,u,b,y)\in T\cap(A\times S_{{\mathbf{A}}}\times B\times S_{{\mathbf{B}}})$, if $y^{\prime}=\delta(u)$ and $u^{\prime}=\delta^{-1}(y)$, then $(u,u^{\prime}),(a,u^{\prime}),(b,y^{\prime}),(y,y^{\prime})$ are Maltsev pairs.
  2. (2)
    For all $(u,u^{\prime})\in\theta$, $(u,u^{\prime})$ and $(u^{\prime},u)$ are Maltsev pairs.
  3. (3)
    $\theta\in\operatorname{Con}({\mathbf{S}}_{{\mathbf{A}}})$.
Proof of Claim 6.10.5.
(1) Suppose $(a,u,b,y)\in T\cap(A\times S_{{\mathbf{A}}}\times B\times S_{{\mathbf{B}}})$ and $y^{\prime}=\delta(u)\in S_{{\mathbf{B}}}$ and $u^{\prime}=\delta^{-1}(y)\in S_{{\mathbf{A}}}$. As $\operatorname{graph}(\delta)\subseteq\operatorname{tr}(T)$ by (6.7), we get $(u^{\prime},u^{\prime},y,y)\in T$. Let $x=d(u,u,u^{\prime})\in S_{{\mathbf{A}}}$. Applying the weak difference term component-wise to $(a,u,b,y)$, $(a,u,b,y)$ and $(u^{\prime},u^{\prime},y,y)$ gives
$(d(a,a,u^{\prime}),x,y,y)\in T.$
By bridge property (B2), we get $d(a,a,u^{\prime})=x$ and thus $(x,y)\in\operatorname{tr}(T)$. As $(x,y)\in S_{{\mathbf{A}}}\times S_{{\mathbf{B}}}$, we then get $(x,y)\in\operatorname{graph}(\delta)$ by (6.7), so $x=u^{\prime}$, which proves $d(a,a,u^{\prime})=d(u,u,u^{\prime})=u^{\prime}$. The other required equalities are proved similarly.
(2) Given $(u,u^{\prime})\in\theta$, pick $(a,b)\in R$ with $\pi_{{\mathbf{A}}}(a)=u$ and $y:=\pi_{{\mathbf{B}}}(b)=\delta(u^{\prime})$. Because $(a,b)\in R$, we then get $(a,u,b,y)\in T$. Thus $(u,u^{\prime})$ is a Maltsev pair by (1). Let $y^{\prime}=\delta(u)$; then $(y,y^{\prime})$ is a Maltsev pair, again by (1). Since $(y,y^{\prime})=(\delta(u^{\prime}),\delta(u))$ and $\delta$ is an isomorphism, it follows that $(u^{\prime},u)$ is a Maltsev pair.
(3) $\theta$ is a subuniverse of $({\mathbf{S}}_{{\mathbf{A}}})^{2}$ by virtue of how $\theta$ is defined, and is reflexive by Claim 6.10.2. It then follows from (2) and Lemma 2.6 that $\theta$ is a congruence. ∎
Claim 6.10.6.
$R$ is the graph of an isomorphism $g:{\mathbf{A}}\cong{\mathbf{B}}$ extending $\delta$.
Proof of Claim 6.10.6.
Suppose $(a,b_{1}),(a,b_{2})\in R$. Let $u:=\pi_{{\mathbf{A}}}(a)$, $y_{i}:=\pi_{{\mathbf{B}}}(b_{i})$, and $u_{i}^{\prime}=\delta^{-1}(y_{i})$ for $i=1,2$. Then $(a,u,b_{i},y_{i})\in T$ and $(u,u_{i}^{\prime})\in\theta$ for $i=1,2$. Hence $(u_{1}^{\prime},u_{2}^{\prime})\in\theta$ by Claim 6.10.5(3), so $d(u_{1}^{\prime},u_{1}^{\prime},u_{2}^{\prime})=u_{2}^{\prime}$ by Claim 6.10.5(2), so $d(y_{1},y_{1},y_{2})=y_{2}$ as $\delta$ is an isomorphism. Also note that $(a,u_{1}^{\prime})$ and $(u,u_{1}^{\prime})$ are Maltsev pairs by Claim 6.10.5(1), so $d(u_{1}^{\prime},a,a)=d(u_{1}^{\prime},u,u)=u_{1}^{\prime}$. Finally, from $(u_{1}^{\prime},y_{1})\in\operatorname{graph}(\delta)\subseteq\operatorname{tr}(T)$ we get $(u_{1}^{\prime},u_{1}^{\prime},y_{1},y_{1})\in T$. Thus applying $d$ to
$(u_{1}^{\prime},u_{1}^{\prime},y_{1},y_{1}),\quad(a,u,b_{1},y_{1}),\quad(a,u,b_{2},y_{2}),$
we get
$(u_{1}^{\prime},u_{1}^{\prime},d(y_{1},b_{1},b_{2}),y_{2})\in T.$
By (B2) we get $d(y_{1},b_{1},b_{2})=y_{2}$ and so $(u_{1}^{\prime},y_{2})\in\operatorname{tr}(T)\cap(S_{{\mathbf{A}}}\times S_{{\mathbf{B}}})=\operatorname{graph}(\delta)$ by (6.7), so $y_{2}=y_{1}$. Thus $y_{1},b_{1},b_{2}$ belong to a common $\kappa$-class and $d(y_{1},b_{1},b_{2})=y_{1}$, which forces $b_{1}=b_{2}$.
This and Claim 6.10.4 prove that $R$ is the graph of a surjective homomorphism $g:{\mathbf{A}}\rightarrow{\mathbf{B}}$. A symmetrical argument proves that $g$ is injective. $g$ extends $\delta$ by Claim 6.10.2(1), which completes the proof of Claim 6.10.6. ∎
Claim 6.10.7.
$R\subseteq\operatorname{tr}(T)$.
Proof of Claim 6.10.7.
Let $(a,b)\in R$, $u=\pi_{{\mathbf{A}}}(a)$ and $y=\pi_{{\mathbf{B}}}(b)$. Then $(a,u,b,y)\in T$. Since $g$ is an isomorphism, we have $g(\mu)=\kappa$, and since $g(a)=b$ and $g(u)=\delta(u)$, we get $(b,\delta(u))=(g(a),g(u))\in\kappa$. Since $\delta(u),y\in b/\kappa\cap S_{{\mathbf{B}}}$, we get $\delta(u)=y$, so $(u,y)\in\operatorname{graph}(\delta)\subseteq\operatorname{tr}(T)$. As $T=\mu\circ T\circ\kappa$, we get $(a,b)\in\operatorname{tr}(T)$. ∎
Now we can finish the proof of Theorem 6.10. Pick any $a\in A\setminus S_{{\mathbf{A}}}$. Let $b=g(a)$, so $(a,b)\in R$. Also let $u=\pi_{{\mathbf{A}}}(a)$ and $y=\pi_{{\mathbf{B}}}(b)$. Then $(a,u,b,y)\in T$, $a\neq u$, and $(a,b)\in\operatorname{tr}(T)$ by Claim 6.10.7. Thus $(a_{1},a_{2},b_{1},b_{2}):=(a,u,b,y)$ satisfies the hypotheses of Claim 6.10.1, and hence by that Claim there exists a bridge $T_{1}$ from $({\mathbf{A}},0,\mu)$ to $({\mathbf{B}},0,\kappa)$ satisfying $T_{1}\subseteq T$, $\operatorname{tr}(T_{1})=\operatorname{tr}(T)$ and (B3), as required. ∎
As a consequence, rooted bridges between meet-irreducible congruences encode similarity between the respective subdirectly irreducible quotients.
Corollary 6.11.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are finite algebras in a locally finite Taylor variety, $\rho\in\operatorname{Con}({\mathbf{A}})$ and $\sigma\in\operatorname{Con}({\mathbf{B}})$ where both $\rho$ and $\sigma$ are meet-irreducible, and $\rho^{+},\sigma^{+}$ are their respective unique upper covers. Then the following are equivalent:
  1. (1)
    There exists a rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$.
  2. (2)
    There exists a bridge from $({\mathbf{A}},\rho,\rho^{+})$ to $({\mathbf{B}},\sigma,\sigma^{+})$ which satisfies (B3).
  3. (3)
    ${\mathbf{A}}/\rho$ and ${\mathbf{B}}/\sigma$ are similar; i.e., ${\mathbf{A}}/\rho\sim{\mathbf{B}}/\sigma$.
Proof.
Let $\overline{{\mathbf{A}}}={\mathbf{A}}/\rho$ and $\overline{{\mathbf{B}}}={\mathbf{B}}/\sigma$. Let $\mu$ and $\kappa$ be the monoliths of $\overline{{\mathbf{A}}}$ and $\overline{{\mathbf{B}}}$ respectively. By the discussion following Definition 4.3, (2) is equivalent to
  1. (2)
    There exists a bridge from $(\overline{{\mathbf{A}}},0_{\overline{A}},\mu)$ to $(\overline{{\mathbf{B}}},0_{\overline{B}},\kappa)$ which satisfies (B3),
which in turn is equivalent to
  1. (2′′)
    There exists a proper bridge from $\overline{{\mathbf{A}}}$ to $\overline{{\mathbf{B}}}$.
Thus (2) $\Leftrightarrow$ (3) by Theorem 3.8, and clearly (2) $\Rightarrow$ (1).
It remains to prove (1) $\Rightarrow$ (2). Assume that $T$ is a rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$. If either $\rho^{+}/\rho$ or $\sigma^{+}/\sigma$ is nonabelian, then both are nonabelian and $\overline{{\mathbf{A}}}\cong\overline{{\mathbf{B}}}$ by Lemma 6.8(2); hence $\overline{{\mathbf{A}}}\sim\overline{{\mathbf{B}}}$, proving (3) and hence (2) in this case.
In the remaining case, both $\rho^{+}/\rho$ and $\sigma^{+}/\sigma$ are abelian. Then $\operatorname{Cov}(\rho)=\{\rho^{+}\}$ and $\operatorname{Cov}(\sigma)=\{\sigma^{+}\}$ by Lemma 6.2(3). Since $T$ is rooted, it contains a bridge from $({\mathbf{A}},\rho,\rho^{+})$ to $({\mathbf{B}},\sigma,\sigma^{+})$. Then by Theorem 6.10, there exists a bridge $T^{\prime}$ from $({\mathbf{A}},\rho,\rho^{+})$ to $({\mathbf{B}},\sigma,\sigma^{+})$ which satisfies (B3), proving (2) in this case as well. ∎
As a special case, we get the following characterization of the “there exists a bridge” relation between irreducible congruences of finite Taylor algebras.
Corollary 6.12.
Suppose ${\mathbf{A}},{\mathbf{B}}$ are finite algebras in a locally finite Taylor variety, and $\rho\in\operatorname{Con}({\mathbf{A}})$ and $\sigma\in\operatorname{Con}({\mathbf{B}})$ where both $\rho$ and $\sigma$ are irreducible. Then the following are equivalent:
  1. (1)
    There exists a bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$.
  2. (2)
    ${\mathbf{A}}/\rho\sim{\mathbf{B}}/\sigma$.
Proof.
As noted following Definition 5.2, every bridge between irreducible congruences is rooted. ∎
7. Conclusion
This paper is the second of three which aim to understand Zhuk’s bridges and their application in algebraic terms. In the first paper [16], we developed the similarity relation between subdirectly irreducible algebras in varieties with a weak difference term. We showed in [16] that, when the monoliths are abelian, the monolith blocks of similar finite subdirectly irreducible algebras naturally inherit the structure of vector spaces over a common finite field. In the current paper, we have connected Zhuk’s bridges to similarity. In particular, if ${\mathbf{A}}$ and ${\mathbf{B}}$ are finite algebras in a Taylor variety and $\rho,\sigma$ are meet-irreducible congruences of ${\mathbf{A}},{\mathbf{B}}$ respectively, then the existence of a rooted bridge from $({\mathbf{A}},\rho)$ to $({\mathbf{B}},\sigma)$ is equivalent to ${\mathbf{A}}/\rho$ and ${\mathbf{B}}/\sigma$ being similar.
The remaining task in our project is to apply the tools of similarity developed in [16] to the bridges Zhuk [18] showed arise from rectangular critical relations (among finite algebras in locally finite idempotent Taylor varieties). In a third paper [15], we will do just that, harmonizing Zhuk’s analysis with the corresponding analysis of Kearnes and Szendrei [11] in the congruence modular case. In particular, we will show that every finite rectangular critical subdirect product $R\leq_{sd}{\mathbf{A}}_{1}\times\cdots\times{\mathbf{A}}_{n}$ in a Taylor variety is described locally by a single linear equation over the finite field arising from the similarity relation.